### v1

# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

Abstract
The sixth generation and beyond communication systems are expected to enable communications of a massive number of machine-type devices. The traffic generated by some of these devices will significantly deviate from those in conventional communication scenarios. For instance, for applications where a massive number of cheap sensors communicate with a base station (BS), the devices will only be sporadically active and there will be no coordination among them or with the BS. For such systems requiring massive random access solutions, a new paradigm called unsourced random access (URA) has been proposed. In URA, all the users employ the same codebook and there is no user identity during the data transmission phase. The destination is only interested in the list of messages being sent from the set of active users. In this survey, we provide a comprehensive overview of existing URA solutions with an emphasis on the state-of-the-art, covering both algorithmic and information-theoretic aspects. Moreover, we provide future research directions and challenges, and describe some potential methods of addressing them.
Index Terms:
Multiple access, random access techniques, grant-free, unsourced random access, machine-type communications, Internet of things
I Introduction
Machine Type Communications (MTC) refers to the communication paradigm tailored for the interaction between devices without direct human intervention. It is a key aspect of the Internet of things (IoT) paradigm, enabling various machines, sensors, and devices to exchange information [1, 3, 2]. MTC plays a crucial role in many applications such as smart cities, industrial automation, and healthcare [5, 1, 4]. MTC exhibits distinctions from human-type communication in several aspects. One notable feature is the ability of MTC to handle a large number of devices with sporadic activity for each device [8, 5, 7, 6]. Specifically, MTC is anticipated to accommodate a density of up to 1 Million devices per square kilometer, and these devices typically operate with low computational and storage capabilities, low duty cycles, and small payloads [5, 6].
Multiple access techniques provide the opportunity for different users to share common communication media simultaneously, thus playing a fundamental role in wireless communication systems [9]. Centralized access (referred to as coordinated access) and random access (also known as initial access and uncoordinated access) are two main categories of multiple access technology [6, 8, 11, 10, 4, 12, 9]. Centralized access refers to a communication protocol or technique where access to a shared communication channel is controlled and managed by a central authority, which allocates distinct communication resources to each user and allows them to transfer data. On the contrary, in random access, users send their messages simultaneously as soon as the new data is generated, without the need for explicit coordination with the central unit [11, 10, 13, 4, 14, 12].
Coordinated multiple access stands as the predominant multiple access mode in the fifth generation (5G) and preceding technologies. Several fundamental coordinated access techniques have been introduced from the first generation (1G) to the fourth generation (4G), including time-division multiple access (TDMA), frequency-division multiple access (FDMA), orthogonal frequency-division multiplexing (OFDMA), code-division multiple access (CDMA), and space-division multiple access (SDMA). These techniques belong to the large category of orthogonal multiple access (OMA) schemes. In the context of 5G networks, the utilization of non-orthogonal multiple access (NOMA) improves the capacity region compared to the conventional techniques [15]. To manage the communication of available users in a system, the mentioned techniques require appropriate scheduling protocols, which are efficient only in the case of few active users [16, 17, 12]. However, to accommodate a large number of users in massive MTC (in beyond 5G and the sixth generation (6G) applications), the coordinated protocols incur excessively high access latency and prohibitive signaling overhead [14, 4, 18, 16, 19, 17, 12].
In contrast to the conventional centralized access, where the base station (BS) allocates resources to the users (devices), users transmit their data without any coordination in the random access setup. Removing the need for scheduling results in some benefits, such as reducing the latency and signaling overhead, which makes the random access scheme an interesting solution for massive MTC systems [17, 14, 4, 18, 16, 19]. ALOHA and grant-free multiple access are two common random access schemes [17, 20].
A widely used random access scheme is the ALOHA multiple access protocol, which enables users to transmit data without prior coordination. In the original description given in [21], Abramson explained that in an ALOHA network, a single broadcast channel is used by multiple communication devices, each of which independently sends packets without coordinating with other devices or following a specific schedule. Specifically, a device sends the entire packet at a randomly selected time, and then waits for an acknowledgment for a period of time. If no acknowledgment is received within the timeout period, it is assumed that a collision has occurred with a packet sent by another device. In such cases, the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occurs or until the process is terminated by the user console. Slotted ALOHA (SA), a variation of ALOHA, is presented as a method for transmitting packets in designated time slots, which doubles the maximum throughput compared to the pure ALOHA protocol [22]. However, there are still some challenges in SA. For instance, in this protocol, several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data [23, 21, 24]. Many studies have focused on improving the effectiveness of SA through various methods such as multiple packet transmission, interference cancellation, and error correction codes [26, 25, 27, 28]. Further analyses of ALOHA-based techniques are provided in [29, 30, 31, 32].
Grant-free random access is another common random access method [5, 12]. In contrast to ALOHA, where users solely transmit data, the grant-free method involves each user transmitting both a preamble and the data sequence [20]. Preamble transmission aims to facilitate collision reduction at the physical layer. Hence, unlike ALOHA, which experiences a significant decline in performance due to collisions, grant-free random access solutions are more resilient to collisions by employing techniques such as user scheduling via a small level of feedback [33], or device-specific patterns such as codebook structures and spreading sequences [17]. So, when there is a large number of users, grant-free random access is a more effective solution than ALOHA. Sourced and unsourced random access schemes are the main categories of grant-free random access. In the first one, both the messages and the identities of users matter to the BS. Therefore, a distinct pilot/preamble is defined for each user, leading to the requirement of a codebook with dimensions proportional to the square of the potential user load. This is impractical in many next-generation wireless network applications with hundreds of millions of potentially connected devices [34, 35].
The so-called unsourced random access (URA), which was introduced by Polyanskiy in [36], by employing a common codebook among users, decouples the user identification and data transmission in the multiple access problems. So, in URA, the receiver is interested in the list of transmitted messages with the decoding done up to a permutation of the messages, and the identity of the users is not important at the first stage. Removing user identity in the URA system enhances its efficiency compared to the sourced random access, enabling it to support a significantly larger number of users [34, 35]. In this paradigm, the BS is connected to millions of cheap devices, a small fraction being active at any given time. The users share all the resources without any priority among them, and they share the same codebook to transmit their payload. Typically, the per-user probability of error (PUPE) is adopted as the performance criterion [36]. The URA paradigm is a promising solution to some next-generation communication applications such as massive machine-to-machine (M2M) communications (one of the main enablers of massive IoT). A block diagram of different multiple access techniques is depicted in Fig. 1.
Following Polyanskiy’s pioneering work, several low-complexity coding solutions are developed in the subsequent literature. In those solutions, by employing successive interference cancellation (SIC) techniques, collisions are resolved up to a certain level depending on user transmit powers. Many techniques, such as random spreading [37], multistage orthogonal pilots [38], and frame slotting [39], have been employed to mitigate the collisions and interference. While a single-antenna receiver in a Gaussian multiple access channel (GMAC) setup is considered in the earlier works, multiple input multiple output (MIMO) receivers, especially massive MIMO, are considered with fading channel in the subsequent literature to support an even greater number of users [37, 38, 39]. Massive MIMO systems [40] help URA by providing high spectral efficiencies, high energy efficiencies, high data rates, and spatial multiplexing gains. This is achieved by generating a substantial number of spatial degrees of freedom (DoF).
Figure 1: Taxonomy of different multiple access techniques.
In the following sections, we provide an overview of the related URA literature, with a focus on GMAC, fading multiple access, and MIMO fading multiple access scenarios. From another aspect, we study the works that put forth coding schemes and those that offer fundamental limit calculations for the URA setup.
The remainder of the survey is organized as follows. In Section II, a general discussion of multiple-access techniques is provided. We start with coordinated multiple-access techniques, both grant-free and grant-based, and find our way to the uncoordinated ones. In Section III, we focus on URA as a new paradigm for uncoordinated multiple-access communications with a massive number of potential devices, enumerate its advantages, and provide some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the fading works with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions.
The acronyms used in this paper are alphabetically summarized in Table I.
Table I: List of acronyms
Acronym Definition
5G fifth generation
5G-NR 5G new radio
6G sixth generation
AD activity detection
AMP approximate message passing
AOA angle of arrival
AWGN additive white Gaussian noise
BCH Bose–Chaudhuri–Hocquenghem
BP belief propagation
BPSK binary phase shift keying
BS base station
CB-MLD covariance-based maximum likelihood detection
CCS coded/coupled compressive sensing
CDMA code division multiple access
CP cyclic prefix
CRC cyclic redundancy check
CS compressed sensing
CSA coded slotted ALOHA
CSI channel state information
DFT discrete Fourier transform
DOF degree of freedom
FBL finite blocklength
FDMA frequency division multiple access
FEC forward error correction
FFT fast Fourier transform
GMAC Gaussian multiple access channel
gOMP generalized orthogonal matching pursuit
HWI hardware impairment
i.i.d. independent and identically distributed
IDMA interleave-division multiple access
IISD iterative inter-symbol decoder
IoT Internet-of-things
IRSA irregular repetition slotted ALOHA
ISAC integrated sensing and communication
LASSO least absolute shrinkage and selection operator
LDPC low-density parity check
LLR log-likelihood ratio
LMMSE linear minimum mean square error
LS least squares
MAC multiple access channel
Acronym Definition
MIMO multiple-input multiple-output
ML maximum likelihood
MMSE minimum mean square error
mMTC massive machine-type communications
MMV-AMP multiple measurement vector AMP
MRC maximal ratio combining
MSE mean square error
MS-MRA multi-stage set-up with multiple receive antennas
MS-MRA-WOPBE MS-MRA without pilot bits encoding
MSUG multi-stage set-up with user grouping
MSUG-MRA MSUG for multiple receive antennas
MTC machine-type communications
MUSIC multiple signal classification
NNLS non-negative least squares
NOMA non-orthogonal multiple access
NP Neyman-Pearson
ODMA on-off division multiple access
OFDM orthogonal frequency division multiplexing
OMA orthogonal multiple access
OMP orthogonal matching pursuit
PMF probability mass function
PUPE per-user probability of error
QPSK quadrature phase shift keying
RA repeat accumulate
RIS reconfigurable intelligent surface
RS Reed-Solomon
SCLD successive cancellation list decoder
SCMA sparse code multiple access
SDMA space division multiple access
SIC successive interference cancellation
SINR signal-to-interference-plus-noise ratio
SISO soft-input soft-output
SKP sparse Kronecker product
SNR signal-to-noise ratio
SPARC sparse regression code
TDMA time division multiple access
TIN treating interference as noise
UE user equipment
UNISAC unsourced ISAC
URA unsourced random access
II Background: Multiple Access
II-A Coordinated Grant-Based Multiple Access
Multiple access refers to the scenario in which more than one user simultaneously shares the system resources. One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combination) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework are TDMA, FDMA, CDMA, OFDMA, and SDMA, which are all grant-based techniques. In TDMA and FDMA, non-overlapping time slots or frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission. On the other hand, in CDMA, all the users can simultaneously utilize all the time-frequency resources through different (nearly) orthogonal code sequences coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing.
All of the approaches described in the previous paragraph are in the category of OMA, where the number of users is strictly limited due to the orthogonal resource assignment. In order to increase the number of supported users, NOMA techniques are investigated in the literature, where multiple users can utilize the same resources to transmit their data, allowing the system to support more users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
The NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [41], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through SIC. On the other hand, code-domain NOMA is similar to classical CDMA in the sense that the users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [42]. Here, the receiver can employ message-passing algorithms utilizing the code structure to differentiate the users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division multiple access [43], low-density spreading aided orthogonal frequency division multiplexing [44], sparse code multiple access (SCMA) [45], and multiuser shared access [46].
II-B Grant-Free Multiple Access
All of the schemes described in the previous subsection are grant-based. That is, they require scheduling grants and coordination with the BS before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5-3 according to 3GPP studies, as the system performance significantly degrades beyond that threshold [47].
Grant-free multiple access is a well-known candidate to enable communication of a large number of users where the users transmit their data by employing the shared system resources without any prior grant from the BS, which reduces the latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks pre-configured by the BS, which means that there is user identity during the data transmission phase, and the receiver can utilize this information for activity detection or user separation. In this line, grant-free NOMA is the prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that, the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection and data recovery, channel estimation, and synchronization with minimal overhead [12].
Among the two main categories of NOMA, the power-domain NOMA is hard to extend to the grant-free scenario as the successful recovery of the transmitted packets at the receiver heavily depends on the power difference of the users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery. For instance, a grant-free version of SCMA is proposed in [48]. Furthermore, the inherent sparsity of the user activity in grant-free multiple access is utilized in [49, 50, 51, 52] through compressive sensing algorithms for multiuser detection.
II-C Uncoordinated Multiple Access
Another candidate solution to address the drawbacks of coordinated multiple access is uncoordinated multiple access, where the users independently exploit the system resources to transmit their data using the same transmission protocol without any coordination with the BS. This scheme is particularly useful when there are stringent latency requirements or when the user activity is random/sporadic, which renders assigning dedicated system resources ineffective.
The first approach for uncoordinated multiple access is the ALOHA protocol [23] proposed in the 1970s, where the users transmit their packets whenever they have data to send. The data packets are assumed to be lost in the case of a collision and they should be re-transmitted. ALOHA is simple; however, it is inefficient as the system throughput is $T_{e}=Le^{-2L}$ with a maximum of $1/{2e}\approx 18\%$ where $T_{e}$ is the system throughput and $L$ is the system load defined as the ratio of the number of transmitted packets to the number of slots. An improvement to the pure version of ALOHA is slotted ALOHA, in which it is assumed that the time is divided into slots and the users can start the transmission only at the beginning of the time slots. Thus, there are no partial collisions, and a packet is assumed as successfully received if there are no other transmissions in the same slot. In this case, the throughput can be calculated as $T_{e}=Le^{-L}$, with a maximum of $1/{e}\approx 37\%$. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 2, and their throughput with respect to the system load is depicted in Fig. 3.
Figure 2: An illustration of pure and slotted ALOHA. Red boxes show the collided packets and green boxes are the successfully transmitted ones.
Figure 3: Throughput comparison of pure and slotted ALOHA.
Although slotted ALOHA doubles the throughput of pure ALOHA, it is still less than 40 %, which is limited. With the motivation to increase the throughput, several variants of ALOHA are proposed in the recent literature. For instance, in contention resolution diversity slotted ALOHA (CRDSA) [27], two copies of the packets are transmitted in two different slots, and in the case that a packet is successfully recovered in one slot, the effect of its replica in the other slot is subtracted. In this way, the maximum throughput can be increased to 52%. This scheme is improved by irregular repetition slotted ALOHA (IRSA) [26], where the packets are repeated a different number of times rather than twice. With the optimization of the repetition distributions by density evolution techniques, IRSA achieves $T_{e}\simeq 0.8$ in practical implementations. An extension of IRSA exploiting forward error correction is called coded slotted ALOHA (CSA) [25], which improves the efficiency even further, approaching a throughput of almost 1 when sufficiently low coding rates are employed. IRSA and CSA are further studied in [29, 30, 31, 32]. Some works also investigate the downlink transmission in an uncoordinated random access setup. For instance, in [53], the downlink of a massive random access setup is considered where the users recover their intended messages from a common message sent by the BS. It is assumed that the BS knows which subset of users are active, however, it is not a valid assumption in an unsourced random access setup.
Figure 4: An illustration of the sporadic user activity.
III URA: A Recent Paradigm
In some massive MTC (or M2M) applications, the total number of devices can be in the order of millions, and they are only active very sporadically as illustrated in Fig. 4. This makes achieving any level of coordination infeasible. As a result, none of the multiple access techniques summarized in the previous section is suitable for enabling the communication of these devices. Specifically, the complexity of the grant-free multiple access is prohibitive due to the codebook size when each potential user is assigned a unique preamble. On the other hand, ALOHA leads to excessive collisions, significantly degrading communication efficiency.
To address this problem, Polyanskiy proposed the URA paradigm in [36] where the data transmission and user identification problems are decoupled via the utilization of a common codebook by all users. So, the receiver is not interested in user identities during the data transmission, with the decoding being done only up to a permutation of the transmitted messages.
Considering a URA model with $M_{r}$ receive antennas, in which $K_{a}$ out of $K_{T}$ users are active at a given frame of length $n$, the received signal matrix in the absence of synchronization errors is written as [38]
$\mathbf{Y}=\sum_{i=1}^{K_{a}}\mathbf{h}_{i}{\mathbf{x}_{i}(\mathbf{u}_{i})+{\mathbf{Z}}}\in\mathbb{C}^{M_{r}\times n},$ (1)
with
$\left\{\begin{matrix}\mathbf{h}_{i}=1,\ \ M_{r}=1\ \ \ \ \ \ \ \ \ \ \ \ \ &\text{GMAC\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\\ \mathbf{h}_{i}\sim\mathcal{CN}(0,1),\ \ M_{r}=1&\text{Fading channel \ \ \ \ \ \ \ \ \ }\\ \mathbf{h}_{i}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M}),\ \ M_{r}>1&\text{MIMO fading channel}\\ \end{matrix}\right.$ (2)
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the length $n$ signal corresponding to the message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of the $i$-th user, $B$ denotes the number of information bits, and elements of the additive Gaussian noise ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ are drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$. Each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine $\mathbf{u}_{i}\ \forall i=1,2,...,K_{a}$ by analyzing the received signal $\mathbf{Y}$.
The existing literature on URA is reviewed in the following sections. However, to address specific constraints within the URA paradigm, we first summarize certain challenges that need to be overcome; namely, the development of bespoke multiuser coding schemes, mitigation of the heavy interference due to the random access nature and the high number of active users, balancing the computational complexity with system performance, and the reduction of preamble/pilot collisions.
III-A Development of Multiuser Coding Schemes
For encoding different users’ messages given the signal model in (1), the most immediate solution is for the active users to map their bit sequences to the rows of a $2^{B}\times n$ codebook in order to select their transmitted signals. The receiver can separate the signals of different users and detect each using detectors such as maximum likelihood (ML) [36]. Nevertheless, it is crucial to note that this coding scheme is only practical for lower values of message length $B$ since its computational complexity is exponential in $B$.
More computationally efficient approaches based on coded compressed sensing and tensor-based techniques (which will be discussed in the following sections) are adopted in [55, 54, 56, 58, 59, 60, 62, 63, 57, 61]. They involve dividing each user’s bit sequence into distinct segments, allowing each segment to be mapped to a codebook with reduced dimensions. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In some other URA coding schemes, each user can encode its entire message sequence (without partitioning it) with strong single-user forward error correction (FEC) codes such as polar or low-density parity-check (LDPC) codes. At the receiver side, estimation techniques such as minimum mean square error (MMSE) are employed to separate the received signals of different users, followed by single-user channel decoding [64, 65, 66, 67, 71, 68, 69, 70, 72, 73, 74, 75, 76, 37, 77, 78, 39, 79, 80, 81, 82, 38]. These algorithms are considered efficient since they manage to strike a balance, avoiding both the high computational complexity associated with ML-based coding schemes and the low accuracy observed in the schemes that segment bit sequences.
III-B Dealing with Heavy Interference
Interference stands out as the primary challenge in multiple access channels. Failure to incorporate strategies for reducing or canceling it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we study the performance of the treating interference as noise (TIN) scheme in a MIMO fading setup, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. Approximating $\frac{1}{M_{r}}\|\mathbf{h}_{i}\|^{2}\approx 1\ \forall i$ in (1), the signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy can be written as
$\alpha_{i}\approx\dfrac{P}{\sigma_{z}^{2}+{(K_{a}-1)}P},$ (3)
where $P$ is the transmitted power of each user. For large enough active user load $K_{a}$, the SINR can be written as
$\alpha_{i}\approx\dfrac{1}{K_{a}}.$ (4)
This equation illustrates that when the number of users surpasses a specific threshold, the SINR becomes inversely proportional with $K_{a}$, and boosting the transmit power does not result in an enhanced SINR. Consequently, a significant rise in $K_{a}$ (equivalent to an increase in interference) adversely impacts the system’s performance. Furthermore, one can observe the huge performance gap between TIN and the SIC-added version of it (TIN-SIC) via the results in [83], where the TIN scheme is compared with TIN-SIC as a strategy designed to mitigate interference.
There are some common techniques in the URA literature to mitigate the interference:
  1. 1.
    Slotted transmission: Dividing the time frame into multiple slots serves to diminish interference by reducing the number of users in each slot [83, 84, 82, 38]. Additionally, slotted transmission results in a shorter duration for the transmitted signal and a reduction in the number of interfering users per slot. Consequently, the overall computational complexity of the URA system is decreased since an increase in these two parameters superlinearly increases the computational workload of the multiuser decoder [39, 38].
  2. 2.
    Sending some preambles/pilots: In some URA schemes, each user randomly selects a preamble/pilot, and transmits it along with its main signal [39, 37, 76, 66, 67, 81, 82, 38]. Upon detecting these transmitted preambles/pilots at the receiver, there is an opportunity to signify the influence of a desired user while mitigating the impact of other interfering ones. For instance, in certain MIMO fading URA schemes, each user selects a pilot and appends it to their codeword before transmission. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. With these estimated channel vectors for various users, interference can be significantly reduced by employing estimators such as MMSE and maximum ratio combining (MRC) [37, 76, 82, 38, 39].
  3. 3.
    SIC: Successive interference cancellation (SIC) is a useful technique for interference mitigation in wireless multiuser scenarios. The idea behind it is to re-encode the successfully detected messages and remove their effects from the received signal to decrease the interference [85, 86, 81, 82, 38]. Note that in fading channels with unknown CSI, the overall performance can be improved with a simple additional step where the corresponding channel coefficients for the user of a detected message are re-estimated before proceeding with the subtraction process [82, 38].
  4. 4.
    Power diversity: An alternative approach for user separation involves randomly distributing users into different groups with varying power levels and adopting an appropriate technique to ensure that the decoding process for different groups is independent of each other [81, 38]. This approach serves to reduce interference within each group. As previously discussed in the context of slotted transmissions, dividing users into distinct groups also leads to a decrease in the computational complexity of the URA system.
III-C Addressing Computational Complexity
To increase the number of active users supported by a URA system, larger values of parameters such as frame length and the number of antenna elements in MIMO systems, are necessary [36, 58, 87]. Additionally, the computational complexity of a URA system escalates with these parameters and the number of active users, resulting in a superlinear dependence of computational load on the active user count in most URA solutions [82, 39]. Consequently, a URA system with a substantial number of active users is anticipated to exhibit a high computational complexity. To mitigate this complexity, various strategies are employed, such as slotted transmissions, and randomly dividing users into different independent groups [82, 39].
III-D Reducing Pilot Collisions
Many existing fading URA solutions employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users randomly and independently pick a preamble/pilot from a common codebook based on a part of their message bits (pilot bits). Hence, if the pilot bits of multiple users are the same (i.e., pilot contamination), a collision occurs and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the collision probability lower than a desired threshold [38] or develop techniques for its mitigation [82].
IV URA over GMAC
IV-A Achievability Bounds
In [36], Polyanskiy introduces the URA paradigm and obtains an achievability bound on its performance on the following GMAC model:
$\mathbf{y}=\mathbf{x}_{1}+\dots+\mathbf{x}_{K_{a}}+\mathbf{z},$ (5)
where $K_{a}$ is the number of active users (i.e., the active user load), $\mathbf{x}_{i}$ is the transmitted signal of the $i$-th user, $\mathbf{y}$ is the received signal and $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is the AWGN noise vector with independent and identically distributed (i.i.d.) elements. Assuming $\norm{\mathbf{x}_{i}}^{2}\leq nP$ is satisfied, the energy-per-bit becomes (actually, it is upper bounded by)
$\frac{E_{b}}{N_{0}}=\frac{nP}{2\log_{2}M},$ (6)
where $M$ is the cardinality of the set of the possible messages $[M]$ and $n$ is the length of the transmission frame. In [36], an $(M,n,\epsilon)$ random access code is defined as follows
Definition 1.
An $(M,n,\epsilon)$ random access code for the $K_{a}$-user channel $P_{Y\mid X^{K_{a}}}$ is a pair of (possibly randomized) maps, the encoder $f:[M]\rightarrow\mathcal{X}^{n}$ and the decoder $g:\mathcal{Y}^{n}\rightarrow\binom{[M]}{K_{a}}$, satisfying
$\frac{1}{K_{a}}\sum\limits_{j=1}^{K_{a}}\mathbb{P}[E_{j}]\leq\epsilon,$ (7)
where $E_{j}\triangleq\{W_{j}\notin\mathcal{L}\}\cup\{W_{j}=W_{i}\hskip 5.69054pt\text{for some }i\neq j\}$ is the error event for the $j$-th user, $W_{1},\dots,W_{K_{a}}$ are independent and uniform on $[M]$ messages, and $\mathcal{L}$ is the list of the estimated user messages that is produced by the decoder. Note that this error definition consists of the misdetection error and the error resulting from users having the same message sequence. Firstly, it does not take into account the false alarm probability. The reason is that since the number of active users is assumed to be known and a fixed-size decoding list of length $K_{a}$ is used at the receiver in [36], the false alarm probability becomes the same as the misdetection error, so there is no need for its inclusion in the error definition. However, the number of active users is not known by the receiver in practice, and needs to be estimated as is done in other URA solutions, either explicitly or implicitly. Therefore, for a more complete analysis of URA systems, both misdetection and false alarm errors need to be taken into account in the error definition. Another point is that it is assumed in this error definition that any instance of two users having the same message leads to an error, which is not valid. However, this assumption does not violate the achievability of the proposed bound as it assumes a higher error.
The achievability bound obtained in [36], which is given below, has been used as a benchmark for the energy efficiency of URA over a real-valued AWGN channel.
Theorem 1 (Random coding achievability [36]).
Fix $P^{\prime}<P$. There exists an $(M,n,\epsilon)$ random-access code for $K_{a}$-user GMAC satisfying power-constraint $P$ and
$\epsilon\leq\sum\limits_{i=1}^{K_{a}}\frac{t}{K_{a}}min(p_{t},q_{t})+p_{0},$ (8)
$\blacksquare$
where $p_{0}=\frac{\binom{Ka}{2}}{M}+K_{a}\mathbb{P}\Bigg{[}\frac{1}{n}\sum\limits_{j=1}^{n}Z_{j}^{2}>\frac{P}{P^{\prime}}\Bigg{]}$, $p_{t}=e^{-nE(t)}$, $E(t)=\max\limits_{0\leq\rho,\rho_{1}\leq 1}-\rho\rho_{1}tR_{1}-\rho_{1}R_{2}+E_{0}(\rho,\rho_{1})$, $E_{0}=\rho_{1}a+\frac{1}{2}\log(1-2b\rho_{1})$, $a=\frac{\rho}{2}\log(1+2P^{\prime}t\lambda)+\frac{1}{2}\log(1+2P^{\prime}t\mu)$, $b=\rho\lambda-\frac{\mu}{1+2P^{\prime}t\mu}$, $\mu=\frac{\rho\lambda}{1+2P^{\prime}t\lambda}$, $\lambda=\frac{P^{\prime}t-1+\sqrt{D}}{4(1+\rho_{1}\rho)P^{\prime}t}$, $D=(P^{\prime}t-1)^{2}+4P^{\prime}t\frac{1+\rho\rho_{1}}{1+\rho}$, $R_{1}=\frac{1}{n}\log M-\frac{1}{nt}\log(t!)$, $R_{2}=\frac{1}{n}\log\binom{K_{a}}{t}$, and $q_{t}$ is defined as
$q_{t}=\inf_{\gamma}\mathbb{P}[I_{t}\leq\gamma]+exp(n(tR_{1}+R_{2})-\gamma).$ (9)
where the random variable $I_{t}$ is defined as
$I_{t}=\min_{S_{0}}i_{t}(c(S_{0});Y|c(S_{0}^{c})$ (10)
where $c(S_{0})\triangleq\sum\limits_{j\in S_{0}}c_{j}$ for $c_{1},\dots,c_{M}\sim\mathcal{N}(0,P^{\prime})$ and the minimum is over all $t$- subsets of $[K_{a}]$. The information density $i_{t}$ is defined as
$i_{t}(a;y|b)=nC_{t}+\frac{\log e}{2}\Big{(}\frac{\norm{y-b}_{2}^{2}}{1+P^{\prime}t}-\norm{y-a-b}_{2}^{2}\Big{)},$ (11)
where $C_{t}=\frac{1}{2}\log(1+P^{\prime}t)$. $\blacksquare$
The idea behind the proof of this theorem is as follows: Let $S=\{W_{1},\dots,W_{K_{a}}\}$ be a random $K_{a}$-subset of $[M]$. We can bound the error event $F_{t}\triangleq\{\lvert S\setminus\hat{S}\rvert=t\}$ with two different methods. The first one is a Gallager-type bound, defining the error events using the generic subsets of size $t$ in $[K_{a}]$ and $[M]\backslash[K_{a}]$, respectively, using the Chernoff bound and applying Gallager’s $\rho$-trick twice. The second one utilizes the information density in (11), considering the event $\tilde{F}=\{I_{t}<\gamma\}$ with the fixed $\gamma$ and $I_{t}$ is given in (10), and taking union over $S_{0}$. The details can be found in [36].
The random coding bound is an information-theoretic performance limit, and it is derived without any complexity constraints as it assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound.
It is assumed in slotted ALOHA that the decoding works if there is no collision and the single-user decoding is successful as the collided packets are discarded in slotted ALOHA protocol. Moreover, as shown in [36], TIN coding satisfies the following approximation
$\log M\approx nC_{TIN}(P)-\sqrt{\frac{nP\log^{2}e}{1+K_{a}P}}Q^{-1}(\epsilon),$ (12)
where $C_{TIN}(P)=\frac{1}{2}\log\left(1+\frac{P}{1+(K_{a}-1)P}\right)$.
Figure 5: Required $E_{b}/N_{0}$ versus the number of active users.
In [36], the performances of these approaches are evaluated assuming that there are $K_{a}$ active users transmitting $k=100$ bits of information through a transmission frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by other GMAC URA solutions for performance analysis, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
The required $E_{b}/N_{0}$ versus number of active users for these schemes are plotted in Fig. 5 along with the random coding bound derived in [36] and a converse bound for $K_{a}$-GMAC from [88]. The results in Fig. 5 show that the conventional random access techniques (i.e., slotted ALOHA and TIN) are energy inefficient for the URA scenario as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. The performance of the grant-based TDMA scheme where the frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame referred to as Orthogonal M.A. (multiple access) is also plotted in the figure (though being infeasible for URA). Note that in all the curves, it is assumed that the single-user codes satisfying the normal approximation in [89] are employed.
In [36], the random coding bound is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors of [90] extend the random coding bound in [36] to the case that the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [91]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [92], where the alarm message needs to be decoded with a lower error rate than the standard ones. Moreover, an extension to the case of asynchronous GMAC is provided in [93], where the delay in transmission can be up to a small portion of the frame. However, the provided achievability bound is a loose bound for the asynchronous case as the authors have employed a fixed time-frame decoder in which all the packets that have some overlaps with the end of the frame are deemed undecodable, leading to huge performance degradation.
IV-B Existing Solutions for GMAC URA
As indicated in the previous subsection, it is shown in [36] that conventional random access techniques are energy inefficient for the URA setup. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems.
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [94], where a concatenated coding scheme is considered combined with the $T$-fold ALOHA protocol. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in the slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [94], each user randomly chooses one slot to transmit its message. The concatenated coding scheme consists of an inner binary linear code which is used to decode the modulo-2 sum of the codewords transmitted within a slot and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [95], where the authors employ a concatenated coding scheme with polar and BCH codes as the inner and outer codes, respectively, and propose a decoding approach exchanging the information between inner and outer decoding.
Figure 6: The slotted transmission structure in [64] based on $T$-fold IRSA.
The authors of [64] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as T-fold IRSA. They propose to split the message into two parts and employ a coding scheme based on a combination of compressed sensing (CS) and LDPC coding at the slot level. Namely, the users pick a preamble determined by the first part of the message from a common codebook, and the second part is encoded by an LDPC code and interleaved before the transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by the non-negative least squares (NNLS) algorithm and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. The general transmission structure in CCS-based URA solutions is depicted in Fig. 6.
$T$-fold IRSA is also utilized in [71] by replacing the LDPC codes with the polar codes and utilizing joint successive cancellation decoding, which provides a significant performance gain over [64]. In [57], the authors employ random signatures to convert the equal-gain channel into a more favorable i.i.d. fading scenario along with trellis-based codes in a $T$-fold ALOHA structure. Their proposed approach provides a low-complexity solution with moderate performance. Moreover, the authors of [96] employ finite-field coding in a slotted structure. More specifically, the users encode their data using a non-binary LDPC code and transmit it after repetition, zero-padding, and interleaving. At the receiver, intra-slot decoding is performed by a joint BP algorithm over a non-binary factor graph. Their proposed approach improves the performance of the slotted schemes in high user loads, compared to its previous counterparts.
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [55], [97, 98, 99, 100, 101, 102, 103], where the messages are divided into sub-blocks and each block is first encoded by an outer code by appending redundant parity bits before being mapped to a CS codebook for the inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer tree code. Based on this main idea, several variations are examined in the literature. Specifically, in [55], the sub-blocks are first encoded by a binary linear code and then mapped to the columns of a common sensing matrix for inner encoding. In [98], sparse regression codes (SPARCs) are exploited as inner codes with approximate message passing (AMP) decoding at the receiver. The authors of [99] improve the performance of the proposed scheme in [98] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [100], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is employed for inner decoding, with the addition of tree decoding. Furthermore, the authors of [101] adapt constant-weight codes as the outer code with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is employed as the outer code in [102]. Finally, in a recent work [103], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the remaining solutions for $K_{a}\leq 225$. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 7.
Figure 7: Encoding-decoding structure in coded compressed sensing (CCS).
Figure 8: A typical decoding structure in random spreading.
Figure 9: Encoding process of sparse IDMA [67].
In another design strategy for URA over GMAC, there is no slotting and the active users utilize the whole transmission frame. In some works following this strategy, the idea is to spread the channel-coded bits through the transmission frame using random signatures [68, 69]. The message is divided into two parts: the first part determines the signature sequence while the second part is encoded by a polar code in [68]. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading based solutions is illustrated in Fig. 8.
In [104], the authors modify the scheme in [68] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. Moreover, [81] improves the performance of the scheme in [68] by letting the users to randomly select a power level from a pre-defined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher received power can be decoded first, and the ones with lower received power are decoded after SIC. This results in a performance improvement over [68] for $K_{a}\geq 150$. The authors of [69] replace the polar codes in [68] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach improves the decoding performance since the information about the partially decoded codewords can be utilized, and it provides the best performance among the schemes based on random spreading.
In random spreading based solutions, each active user occupies the entire frame by random spreading of channel-coded bits for data transmission. In some other works, the active users distribute their codeword bits onto the transmission frame in different ways. For instance, in the sparse interleave-division multiple access (IDMA) approach [67], the message bits and the channel frame are divided into parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The rest of the bits are encoded by an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted along the rest of the channel frame. This encoding process is illustrated in Fig. 9. At the receiver, after decoding the preamble part via the least absolute shrinkage and selection operator (LASSO) technique, a single joint Tanner graph is utilized for message recovery, exploiting the information of the location of transmitted symbols from the identified preamble.
A similar approach to [67] employing polar codes is adopted in [105], however, different code lengths are used among the users and TIN is utilized as the decoding algorithm. Moreover, in [106], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts with the first part being mapped to a column of a common sensing matrix which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part followed by a multiuser decoding algorithm proposed in the second part. Furthermore, the authors of [70] propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components. Furthermore, enhanced spread spectrum ALOHA (E-SSA) [107], an asynchronous random access protocol combining ALOHA with direct sequence spread spectrum is adapted for URA over GMAC in [108]. The authors employ a wrap-around version of E-SSA to be able to compare their proposed solution with the existing synchronous solutions and benchmarks.
In a recent work [109], the authors employ on-off division multiple access (ODMA) [110] in the context of URA, where the active users exploit a small portion of the transmission frame to transmit their data encoded by a repeat-accumulate (RA) code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors of [111] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [112], which improves the scheme in [109] by employing a non-binary LDPC code and resolving the collisions of two users.
A summary of the available coding schemes in URA over GMAC can be found in Table II, accompanied by a comparison of their performance in terms of energy efficiency in Fig. 10 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. The results in Fig. 10 demonstrate that random spreading based schemes (Pradhan et al., 2020, Ahmadi et al. 2021, Pradhan et al., 2021) offer good performance for a lower number of active users; however, their performance degrades with the increasing number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA (Yan et al., 2023, Ozates et al., 2024, Yan et al., 2024) outperforms the other approaches except the solution proposed in Nassaji et al., 2024 in the high multiuser interference regime due to its sparsity, while it performs well in the low multiuser regime when combined with polar coding. Moreover, CCS can exhibit very good performance for a high number of active users when the rate loss due to the redundant parity bits is alleviated as shown in Nassaji et al., 2024. Regarding the comparison with the achievability bound, the proposed schemes can perform similar or slightly better for $K_{a}\leq 100$, and the gap is a maximum of 0.8 dB for $K_{a}\leq 300$.
Table II: Main coding schemes in URA over GMAC
Paper Contribution and Techniques
Polyanskiy, 2017 [36] information-theoretic formulation of URA. Random coding achievability and converse bounds.
Ordentlich et al., 2017 [94] slotting based scheme with concatenated coding combined with $T$-fold ALOHA
Vem et al., 2019 [64] slotting based scheme with LDPC coding, preamble transmission, and SIC
Marshakov et al., 2019 [71] slotting based scheme with preamble transmission, polar coding, and SIC
Amalladinne et al., 2020 [55] CCS-based scheme with inner CS and outer tree coding.
Amalladinne et al., 2022 [99] CCS-based scheme with the interaction between the inner AMP and the outer BP
Nassaji et al., 2024 [103] CCS-based scheme eliminating the outer tree code.
Pradhan et al., 2022 [67] sparse spreading based scheme with preamble transmission, LDPC coding and SIC
Han et al., 2021 [70] sparse spreading based scheme with convolutional coding and SIC
Pradhan et al., 2020 [68] random spreading based scheme with polar coding and SIC
Ahmadi et al., 2020 [81] random spreading based scheme with power diversity, polar coding, and SIC
Pradhan et al., 2021 [69] random spreading based scheme with LDPC coding, soft cancellation, and SIC
Yan et al., 2023 [109] ODMA based scheme with RA coding and soft cancellation
Ozates et al., 2024 [111] ODMA based scheme with polar coding and SIC
Yan et al., 2024 [112] ODMA based scheme with non-binary LDPC coding and soft cancellation.
Figure 10: Performance comparison of the coding schemes in URA over GMAC for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature.
V URA over Fading Channels
Since GMAC is not a very realistic channel model for wireless communications, a number of studies explore URA over Rayleigh fading channels [67, 83, 84, 114, 116, 115, 113, 56, 72, 74, 75, 73]. In [83, 84, 113, 114, 115], the authors derive some performance bounds for the URA setup over a Rayleigh fading channel. For instance, the authors of [115] present a converse bound based on the results from [117] and meta-converse from [118], stated below.
Theorem 2 (Converse bound [115]).
Let
$L_{n}=n\hskip 2.84526pt\log(1+PG)+\sum\limits_{i=1}^{n}\left(1-\absolutevalue{\sqrt{PG}Z_{i}-\sqrt{1+PG}}^{2}\right),$ (13)
$S_{n}=n\hskip 2.84526pt\log(1+PG)\sum\limits_{i=1}^{n}\left(1-\frac{\absolutevalue{\sqrt{PG}Z_{i}-1}^{2}}{1+PG}\right),$ (14)
where $G=\norm{h}^{2}$ and $Z_{i}\sim\mathcal{CN}(0,1)$. Then, for every $n$ and $10<\epsilon<$ 1, any $(M,n-1,\epsilon)$ code for the quasi-static $K_{a}$ multiple access satisfies
$\log(M)\leq\log(K_{a})+\log\frac{1}{\mathbb{P}(L_{n}\geq n\gamma_{n})},$ (15)
where $\gamma_{n}$ is the solution of $\mathbb{P}\left[S_{n}\leq n\gamma_{n}\right]=\epsilon$.                                             $\blacksquare$
In [84] and [113], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. Authors in [114, 115] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [83], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering slotted structure and perfect SIC in their analysis. Their proposed non-asymptotic achievability bound is tighter than the ones in [84, 113, 114, 115] however, the assumption of perfect SIC may not be realistic.
V-A Single-Antenna Fading Channels
We first take a look at the existing URA fading solutions assuming a single-antenna receiver. The works in [55, 67, 119, 120, 83, 114, 115, 56, 72, 74, 116, 75, 73] introduce practical coding schemes tailored for the URA scheme on fading channels. Within this collection of works, some assume that the received signals from all users are synchronized [67, 55, 83, 114, 115, 56], while others take into account asynchronous transmission [116, 72, 74, 75, 73]. To estimate the delay of asynchronous signals, [116, 72, 74, 73] use OFDM transmission to convert the time delay to the phase shift in the frequency domain, while the authors in [75] consider time-domain transmission where the delay of a signal is detected via correlation-based approach.
Focusing on the coding schemes in the context of Rayleigh fading channels, we have the following classification:
TIN: Similar to GMAC, the TIN solution for fading channels operates under the assumption of the presence of just a single user and treats other interfering users as noise. Consequently, this approach results in an exceptionally low average signal-to-noise ratio (SNR) due to the fact that it encounters interference-plus-noise as the system’s overall noise level, leading to a very low performance.
Figure 11: Encoding (top) and decoding (bottom) structures of a CCS scheme with 4 slots [55].
CCS [56, 75, 120]: As explained in the previous section, CCS was originally proposed for URA over GMAC. Nevertheless, it has been applied to the fading scenario as well in several works. The encoding and decoding structures of a CCS-based scheme with four slots are shown in Fig. 11, where each user’s signal is divided into four parts, with each one being appended to a parity-check bit. The result is mapped to a codebook to construct the pilot of the user in a slot. Also, at the receiver, sub-messages of different users are detected at each slot, and corresponding ones in different slots are stitched together to form the complete message sequence of a user.
In [56], the authors modify the tree decoder of the CCS scheme in [55] by replacing it with two codes capable of correcting $t$ errors: the first one modifies the conventional tree code by giving it the ability to recover $t$ errors (called $t$-tree), and in the second one, Reed-Solomon (RS) code is employed. In [120], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS encoding, with AMP at the receiver. Moreover, in [75], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all potential delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays when performing CS decoding.
Polar coding with SIC [83]: We next delve into the encoding and decoding procedures of the schemes employing polar codes in conjunction with TIN-SIC. The frame in this scheme is divided into $S$ slots. Each user randomly selects a slot to transmit its polar coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by an SIC. This is because, different users in a slot are prone to different fading coefficients, some of them with larger coefficients are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating others as noise, its channel coefficient is estimated using the orthogonal matching pursuit (OMP) technique. Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is demonstrated in Fig. 12.
Figure 12: Encoding and decoding process of the polar code with SIC scheme in [83] in a slot with $K$ users.
There are also some other solutions focusing on URA over fading channels with a single-antenna receiver. For instance, in [119], the active users utilize the entire transmission frame to transmit their data. Specifically, the channel frame is divided into pilot and data parts. The active users pick a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits along the rest of the transmission frame with permutation and scrambling in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits. Furthermore, the authors of [116] study URA over frequency-selective channels and propose OFDM to mitigate the channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
LDPC with joint decoding [114, 115]: For encoding, these algorithms partition the time frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers two significant benefits: it reduces the computational load on the decoder by reducing the length of the LDPC signal, and mitigates per-slot interference by distributing users across slots.
The schemes utilize an iterative belief propagation (BP) decoder, as illustrated in Fig. 13. The graph incorporates four distinct node types: variable nodes (in red), and check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to simultaneously update the estimated values of both the channel coefficients and the symbols for each user.
Figure 13: Iterative BP decoder for a two-users LDPC-based scheme with joint decoding [114].
Figure 14: Performance comparison of the coding schemes in fading URA with a single-antenna receiver, with a target PUPE of $0.1$, $n=30000$, and $B=100$ (the common set of parameters in the literature).
LDPC coding with SIC [72, 73]: This coding design builds upon the schemes presented in [114, 115], with the key distinction of incorporating OFDM techniques to mitigate the effects of the asynchronous transmission by converting the time delay into a phase shift, and employing SIC instead of joint decoding.
The performance of different coding schemes in the single-antenna fading URA is compared in Fig. 14 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The CCS-based scheme in Andreev et al., 2022 exhibits lower energy efficiency compared to the other ones since it employs a sub-optimal coding scheme, while others use strong channel codes such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in Kowshik et al., 2020 is inferior to that of the ones in Pradhan et al., 2022, Kowshik et al., 2019, Andreev et al., 2020, which can be attributed to the absence of an efficient SIC block in this LDPC-based scheme. On the other hand, the TIN-SIC approach in Kowshik et al., 2019, Andreev et al., 2020 offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when it is combined with polar coding as in Andreev et al., 2020. Nevertheless, the TIN-SIC scheme in Andreev et al., 2020 is outperformed by the spreading based scheme in Nassaji et al., 2022, and the ODMA-based scheme employing polar codes in Ozates et al., 2024, offering the best performance among the proposed schemes for single-antenna fading case.
A summary of URA studies over fading channels involving a single receive antenna is given in Table III.
Table III: Summary of available coding schemes in URA over fading channels.
Paper Contribution Description Synchron. vs. Asynch.
Kowshik et al., 2019&2021 [84, 113] Bounds Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron.
Andreev et al., 2020 [83] Bounds & coding scheme: polar code with SIC Non-asymptotic achievability and converse bounds for unknown CSI + a new coding scheme containing slotting, polar code, and SIC. Synchron.
Kowshik et al., 2019&2020 [114, 115] Bounds& LDPC code with joint decoding Achievable and converse bounds for known and unknown CSI in asymptotic and non-asymptotic manners + a new coding scheme containing LDPC code, slotting, and iterative joint decoding algorithm. Synchron.
Andreev et al., 2022 [56] Coding scheme: CCS-based Improving CCS in [55] by replacing the conventional outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron.
Yun et al., 2024 [120] Coding scheme: CCS-based A slotted scheme including CCS and AMP in a block fading model. Synchron.
Nassaji et al., 2022 [119] Coding scheme: Spreading-based A non-slotted scheme including, preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron.
Andreev et al., 2019 [72, 73] Coding scheme: LDPC code with SIC A new coding scheme containing LDPC code, OFDM, slotting, and SIC. Asynch.
Chen et al., 2017 [74] Coding scheme: OFDM with SIC A new coding scheme containing OFDM and SIC. Asynch.
Ozates et al., 2023 [116] Coding scheme: OFDM with SIC A new coding scheme in the frequency-selective channel, containing OFDM, pilot transmission, slotting, polar code, and SIC. Asynch.
Amalladinne et al., 2019 [75] Coding scheme: CCS-based They extend CCS in [55] to accommodate asynchronous scenarios. Asynch.
V-B MIMO Fading Channels
We now turn our focus to the URA solutions for MIMO fading channels. In [87, 121], achievability and converse bounds are obtained for URA over MIMO Rayleigh fading channels, whose ML-based achievability bound is obtained by employing a projection decoder in [122]. Furthermore, several studies have investigated practical coding schemes for the MIMO setting, which can be categorized into four groups: CCS-based, tensor-based, pilot-based, and Bayesian approaches [60, 62, 58, 59, 123, 124, 76, 37, 77, 78, 39, 79, 125, 80, 63, 82, 38], as depicted in Fig. 15.
Figure 15: Coding strategies in URA over MIMO fading channels.
As the computational complexity of mapping bit sequences of each user to a codebook exponentially increases with the message length $B$, CCS-based algorithms are employed in which the bit sequence of each user is divided into $S$ sub-messages [58, 59]. Subsequently, these sub-messages are mapped to pilot codebooks with lower dimensions (see the top part of Fig. 11) to obtain $S$ different pilots. Then, pilots are transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot, and they are then combined (see the lower part of Fig. 11) to form the complete length-$B$ bit sequence for each user. To establish connections between these sub-messages, the authors in [58] incorporate parity-check bits into each sub-message, while [59] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the necessity for additional parity bits enhances the performance of [59] in comparison to [58].
Eliminating the parity bits in [58] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [123, 124, 126, 127]. To stitch the decoded sub-messages, the authors of [123] employ a joint Bayesian decoding exploiting the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [126] and the similarity of the estimated channel vectors are used in [127]. The structure of the inner SPARC code allows parallel single-user outer decoding of device messages in [124], which mitigates the need for the parity check bits. The uncoupled structure is also employed in [128] to design a URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [129], where a long sub-block that is recovered in the first phase is used for channel estimation, and very short lengths of sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [130] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller codebook in both the codeword length and the number of codewords.
Figure 16: Encoding and decoding structures of tensor-based schemes, where each user divides its message sequence into $d$ parts (here $d=2$), then applies Kronecker product on them [62].
Another trend in MIMO URA is the tensor-based schemes, where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor) [60, 61, 62, 63]. For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [60], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit-sequence is detected. This scheme is extended to the case with the presence of timing offsets in [61] by employing OFDM. On the other hand, the authors of [62] propose a block-term decomposition for the decomposition of the 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. The overall coding structure of tensor-based methods is depicted in Fig. 16.
Figure 17: Encoding structure of pilot-based schemes.
Figure 18: Decoding structure of pilot-based schemes.
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems [82, 38, 39, 78, 77, 37, 125, 76, 131], where, similar to the GMAC case, the idea is to divide the bit sequences into pilot and data parts. The transmitted signal of each user also consists of two parts (see Fig. 17 for details): the pilot part which is obtained by mapping the pilot bits to a non-orthogonal codebook [76, 37, 78, 77, 39, 125], or a codebook consisting of multiple stages of orthogonal pilots [38], and the data part which is the encoded by a a polar [38, 76, 37, 39] or an LDPC code [78, 77]. For decoding, as shown in Fig. 18, these schemes iteratively employ the following steps:
  • At the initial stage, they detect the active pilots using different algorithms such as generalized orthogonal matching pursuit (gOMP) [39, 125], AMP [76, 77, 131], energy detector [37], NP hypothesis testing [38] or CB-MLD [78].
  • The channel vectors corresponding to each detected pilot are estimated by techniques such as MMSE [76, 37, 39, 131, 125], MRC [38] or AMP [78, 77].
  • Using the estimated channel coefficients, they obtain soft estimates of the transmitted codewords via MRC [76, 38, 39, 125], MMSE [37] or AMP [78], which are used for generating log-likelihood ratio (LLR) values to be fed a single-user decoder [38, 76, 37, 78, 39, 131] or a joint factor graph [77].
  • The contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the next iterations (this step is not employed by all the pilot-based schemes). Note that in [38, 37, 39, 125], the channel vectors are re-estimated using the successfully decoded codewords to be utilized for SIC.
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is extended to the fading scenario as well for the case of a single-antenna receiver in [111], and a massive MIMO receiver in [125]. The main idea is to allocate a part of the channel frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 19. Note that if the channel is a GMAC, the whole channel frame is used for data transmission.
Figure 19: An illustration of the ODMA idea for the fading scenario with the pilot and data parts. Colored boxes show the utilized time instances in the data part.
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently, hence the error from one block is propagated to the next one. To resolve this issue, there is also a new set of MIMO URA schemes that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal, and their corresponding channel coefficients [80, 79]. Specifically, in [79], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Besides, the authors of [80] propose to encode the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
In Fig. 20, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients. The results in Fig. 20 show that due to adopting strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes (Fengler et al., 2022, Gkagkos et al., 2023, Ahmadi et al., 2024, Ozates et al., 2024) usually outperform the tensor-based (Decurninge et al., 2020) and CCS-based (Fengler et al., 2021) schemes for which sub-optimal coding schemes are utilized, while the SKP coding-based scheme in Han et al., 2023 following the Bayesian approach offers a similar performance to the best performing pilot-based schemes.
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution can be employing RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of the network by enabling line-of-sight paths between the transmitters and the receivers in problematic environments with many blocking obstacles [132, 133]. The authors in [134, 63] show that the use of RIS can improve user connectivity in URA by creating high-quality links between the BS and the users. More specifically, the RIS-based schemes in [134, 63] operate in two phases. In the first phase, called the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, called the data phase, transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [134] outperforms the one in [63], with a power saving of up to 10 dBm. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, and a more suitable metric in the RIS design in [134], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [63]. The authors improve the RIS design metrics of [134] in [135] and obtain some bounds on its performance.
Figure 20: Comparison of URA schemes for MIMO fading channels with a target PUPE of $0.05$, $n=3200$, and $B=100$ (the common set of parameters in the literature).
Conventionally, in MIMO URA systems, the ideal hardware units are assumed at both the BS and the UE side. However, due to the massive number of users and the antenna elements, the hardware needs to be inexpensive, and these elements are prone to HWIs such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWIs is considered in [136], where the solution developed in [39] is adapted to the presence of HWIs. Furthermore, a channel model with one single dominant path between a user and the BS is considered in [137], and an approach similar to the one in [131] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [138], namely, the users can start or end their transmission in any slot and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing.
Table IV: Summary of available coding schemes in URA with MIMO fading channels.
Paper Contributions Description
Gao et al., 2023 [87] Bounds Achievability and converse bounds.
Fengler et al., 2021 [58] Coding scheme: CCS-based The pilots at each slot are found using a newly proposed coordinate-wise activity detector. Then, the detected pilots of different slots are stitched together using parity bits.
Shyianov et al., 2021 [59] Coding scheme: CCS-based Instead of employing parity-check bits (as in [58]), the correlation between different slots is used for stitching pilots, hence, the energy efficiency is improved.
Jiang et al., 2021 [129] Coding scheme: CCS-based A two-stage scheme with a long sub-message in first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage.
Decurninge et al., 2020 [60] Coding scheme: tensor-based Users’ signal separation and detection is done by tensor decomposition.
Luan et al., 2022 [62] Coding scheme: Tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [60].
Shao et al., 2022 [63] Coding scheme: tensor-based Users communicate with the BS through the RIS. For detecting the transmitted signals, a Bayesian approach is employed.
Fengler et al., 2022 [76] Coding scheme: pilot-based A pilot and a polar codeword are transmitted sequentially. For decoding, the active pilots and corresponding channel coefficients are detected using MMV-AMP and MMSE, respectively, and a polar decoder is employed for data detection.
Ahmadi et al., 2024 [38] Coding scheme: pilot-based Each user transmits its multiple stages of orthogonal pilots along with a polar codeword through a randomly selected slot. An energy detector, an MRC, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, an LS-based SIC removes the contribution of successfully decoded signals. Also, an efficient strategy is proposed for randomly dividing the users into different groups.
Gkagkos et al., 2023 [37] Coding scheme: pilot-based A pilot and a random spread polar codeword are transmitted sequentially. An energy detector, an MMSE, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, a SIC removes the contribution of successfully decoded signals.
Su et al., 2023 [77] Coding scheme: pilot-based A pilot is transmitted along with an LDPC coded and modulated codeword. At the receiver side, AMP is used for pilot detection and channel estimation, and an LDPC decoder is used for data detection. The selected pilot specifies the indexes over which every symbol of the LDPC codeword is transmitted.
Su et al., 2023 [78] Coding scheme: pilot-based A pilot and an LDPC codeword are spread and transmitted consecutively. At the receiver’s end, an ML algorithm detects active pilots, an AMP technique estimates the corresponding channels, and a SIC block is applied following the LDPC decoder.
Ozates et al., 2024 [39] Coding scheme: pilot-based The transmission frame is slotted, and each user randomly selects a slot to transmit a pilot along with a polar codeword. For decoding, a generalized OMP and an MMSE estimator are used for pilot detection and channel estimation, the polar decoder is used for data detection, and SIC removes the contribution of the decoded messages.
Ozates et al., 2024 [125] Coding scheme: pilot-based The transmission frame is divided into pilot and data parts, and each user transmits a non-orthogonal pilot followed by its polar codeword based on ODMA. At the receiver, an iterative decoder employing gOMP, LMMSE, MRC, single-user decoding, and SIC is utilized.
Han et al., 2023 [79] Coding scheme: Bayesian approach Employs FEC code, and sparse spreading.
Jiang et al., 2023 [80] Coding scheme: Bayesian approach Combines sparse regression codes for encoding and a message-passing algorithm for decoding.
While most URA works consider the transmission over i.i.d. Rayleigh fading channels, the more practical Rician fading case is considered in [139] through an application of coded compressed sensing. A correlated fading scenario is tackled in [140]. Asynchronous MIMO URA is studied in [141, 143, 142]. A slotted concatenated coding scheme for the asynchronous URA is proposed in [141] based on the SPARCs and tree codes in an OFDM system. The authors of [142] tackle both the timing and frequency offsets by proposing an asynchronous URA scheme by employing SPARC, LDPC codes, and multistage orthogonal pilots. Moreover, symbol asynchronicity in URA is investigated in [143]. Moreover, some works have investigated URA in other communication setups such as cell-free MIMO [144] and erasure [145] channels. For instance, a cell-free setup is considered in [144] where the coverage area is divided into multiple regions each served by a receiver unit. The users in each region employ a common codebook different from other regions to transmit their messages, and the received signals at the receiver units are jointly processed in a central unit.
A summary of MIMO URA schemes is provided in Table IV.
VI Remarks and Future Research Directions
In this survey, we had a comprehensive look at the URA literature with an emphasis on the state-of-the-art. We grouped the different URA solutions and explained the representative work in each in more detail while clarifying its connection to related works. Moreover, we compared the works in the literature both setup-wise and performance-wise in order to have a clearer view of the existing URA solution.
As evident from this survey, there have been numerous works on URA in recent years. However, there are still some issues, some critical, that need to be investigated further. We will next outline some of these issues and provide several future research directions. Moreover, we highlight potential methods of addressing them.
VI-A Asynchronous URA
In URA, the common assumption is that the users transmit simultaneously; or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of the users and devices in envisioned URA applications such as massive sensor networks, such transmission coordination is very challenging. This makes asynchronous URA an important research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. The reason is that due to the massive number of potential devices, and hence diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous.
There have been some attempts to address the asynchronous setup [73, 72, 116, 141, 142, 146, 93], most of which try to convert the problem to one similar to the synchronous case, e.g., by employing a limited decoding time-frame. For instance, in [73], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users. However, it is shown that if the asynchronicity is dealt with properly, it can lead to performance improvement even compared to the synchronous case [147]. One way to realize this is the use of frameworks that are more suitable for the asynchronous case. For instance, using sliding window frameworks with an (almost) infinite time horizon, instead of fixed-time time-frames.
VI-B Robust URA
Looking at the URA literature, we can observe that almost all existing URA schemes are sensitive to the choice of parameters in their proposed algorithm. That is, the proposed algorithms need to be optimized every time a system hyper-parameters, such as active user load or the target PUPE changes, in order to have an acceptable performance. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyper-parameters.
Another robustness aspect is due to the imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments where the employed hardware is not as ideal as assumed, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, which is very probable in many URA applications where the devices are actually cheap and low-precision sensors. Only one work in the URA literature has addressed this issue [136], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensations are statistically modeled as additive and multiplicative noises. Therefore, the analysis of different hardware impairments on the performance of the existing URA solutions and providing robust solutions for them is an interesting research direction.
VI-C Machine Learning-Aided URA
Current solutions look at the URA problem in a single shot. However, the transmission is continuous and spans several frames in many practical applications. In these cases, machine learning techniques such as reinforcement learning can be used to adapt to the environment and learn suitable transmission strategies. Machine learning techniques can also be used to simplify some blocks in the decoding process, such as user activity detection (similar to the deep learning based solutions in [148]) and channel estimation. For instance, in [149], the authors show that employing a dictionary learning based solution obviates the need for pilots for channel estimation in the fading URA scenarios while maintaining an acceptable performance. Machine learning techniques can also be utilized to make the system more robust against changes in the environment or system hyper-parameters such as the active user load, especially as the design parameters of many existing URA solutions heavily depend on these hyper-parameters.
It can be vice versa as well: URA techniques can be employed to better realize some machine learning based solutions. For instance, in federated learning, the receiver is only interested in the sum of the transmitted signals (a function of neural network weights or their updates), not them separately. The users transmit simultaneously in over-the-air federated learning, which exploits the additive nature of the wireless channel. This alongside the huge number of potential users in envisioned federated learning applications, makes URA a suitable framework for the physical layer [150, 151]. Note that in this case, PUPE is not a suitable metric as the receiver is not interested in the transmitted messages themselves but only in their summation. So, new metrics (even higher-level machine learning based ones such as accuracy) need to be employed to better capture the system performance.
VI-D Integrated Sensing and Communication in URA
Existing research on URA has focused primarily on the transmission of information from communication users. However, with the evolution of 6G and beyond, new functionalities such as sensing, computing, and security are becoming critical alongside communication. As the hardware architectures and signal processing methods for these tasks are mostly similar, jointly adopting them significantly reduces hardware costs compared to separate systems while maintaining comparable performance [152]. A promising example of such joint designs is integrated sensing and communication (ISAC) systems.
In [153], the problem of integrating sensing and communication is considered in a URA setup, where a massive number of unsourced and sporadically active users are connected to the system. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievable bound for the UNISAC to show its efficacy in sensing the sensing users as well as in communicating with the communication ones in a URA setup. However, there are still many aspects of an ISAC system in a URA setup that need to be investigated. For instance, the problem becomes more challenging when the sensing users are passive in contrast to the active assumption in [153].
VI-E URA with Feedback
In the URA literature, it is assumed that there is no coordination between the users and the receiver. However, there can be some level of coordination as long as the signal sent from the receiver is not user-specific as there is no user identity in URA. For instance, the receiver can broadcast a common pilot/beacon to the users so that they can estimate their channel coefficients or at least their large-scale components depending on their dedicated time slots. Just note that the time dedicated to this operation needs to be considered as extra overhead when comparing with the solutions without feedback. This issue is partly tackled in [154, 155, 156]. In [154], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques; while in [155], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a pre-defined threshold. One application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [156] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
Recall that the main novelty of URA is providing a unified framework to decouple user identification and data transmission problems. This makes it a fit solution in applications where user identification is not feasible at all due to the massive number of potential users. However, URA is still applicable and a good candidate in non-massive (but with still a considerable number of potential users) scenarios where user identification is still feasible. In these scenarios, conventional authentication and acknowledgment schemes impose a huge overhead on the system. This issue is tackled in [157].
VI-F RIS-Aided URA
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play an important role in improving the overall system performance by providing a substitute communication link with potentially many degrees of freedom. Only a few works have considered the RIS-aided URA scenario [134, 63], which both assume a conventional passive RIS.
For moving forward, a few important points can be mentioned. Firstly, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [158] and metasurface-based RIS [159], in order to find suitable candidates. Secondly, the employment of massive arrays such as extra large massive MIMO and RIS leads to some active users being in the near field of these arrays, which makes hybrid near field-far field analysis a necessity. This is in contrast with the existing URA literature which are all in the far field regime. Thirdly, channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the high number of RIS elements and potentially massive number of antennas at the receiver (as the benefits of a RIS system is amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even active users, which renders most existing channel estimation schemes in RIS infeasible.
VI-G Multi-Class URA
The common assumption in the URA literature is that all the users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, which is called multi-class URA is addressed in only a few works in the literature [160, 161] in the coded compressed sensing framework. Hence, a further investigation of multi-class URA in different frameworks can be interesting.
VI-H Standardization
There has not been any effort in the standardization of URA, neither for incorporating it in the next updates of the current protocol nor for future generations such as 6G. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that preferably can support both the current protocols, perhaps with small modifications, and the high demands of the future ones. For instance, in [162], with a small modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with even a moderate number of active users (10 to 50 users), employing an IDMA-based URA solution.
There are two main challenges in incorporating URA in the current protocols such as 5G-NR. The first one is the small number of available preamble sequences in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second one is the limited number of access patterns that can be realized which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This shows that one important step in the realization of URA in future protocols is to considerably increase the available pilot or transmission patterns.