### v2

# Unsourced Random Access: A Comprehensive Survey

Abstract
Multiple access communication systems enable numerous users to share common communication resources simultaneously, playing a crucial role in wireless networks. With the emergence of the sixth generation (6G) and beyond communication systems, supporting massive machine-type communications with sporadic activity patterns is expected to become a critical challenge. Unsourced random access (URA) has emerged as a promising paradigm to address this challenge by decoupling user identification from data transmission through the use of a common codebook. This survey provides a comprehensive overview of URA solutions, covering both theoretical foundations and practical implementations. We present a systematic classification of URA solutions across three main channel models: Gaussian multiple access channels (GMACs), single-antenna fading, and multiple-input multiple-output (MIMO) fading channels. For each category, we analyze and compare state-of-the-art solutions in terms of performance, complexity, and practical feasibility. Additionally, we discuss critical challenges such as interference management, computational complexity, and synchronization issues. The survey concludes with promising future research directions and potential methods to address existing limitations, providing a roadmap for researchers and practitioners in this rapidly evolving field.
Index Terms:
Multiple access, RA techniques, grant-free, unsourced RA, machine-type communications, Internet of things
I Introduction
The rapid proliferation of Internet of Things (IoT) devices is driving a fundamental shift in wireless communication systems. Current estimates project over 15 billion connected IoT devices by 2023, with the expected increase to a staggering number of over 38 billion devices by 2029 [1]. One example of such systems is machine-type communications (MTC) (also known as machine-to-machine (M2M) communications), which, as a key aspect of the IoT paradigm, enables various machines, sensors, and devices to exchange information without direct human intervention. MTC plays a crucial role in many applications ranging from industrial automation and smart cities to healthcare monitoring and environmental sensing [2]. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer), which typically operate with low computational and storage capabilities, low duty cycles, and small payloads [3]. This massive scale presents unprecedented challenges for wireless access, as traditional communication paradigms were designed for a relatively small number of high-data-rate users rather than an enormous number of devices with sporadic, low-rate transmissions.
Multiple access (MA) techniques, which enable different users to share a common communication medium simultaneously, have been fundamental to wireless communication systems [4]. MA technologies can be broadly categorized into centralized access and random access (RA). In centralized access (also known as coordinated access), a central authority controls and manages access to the shared communication channel by allocating distinct resources to each user for data transmission. This approach has been the predominant MA mode in the fifth generation (5G) and previous cellular technologies. However, centralized coordination becomes impractical for massive IoT scenarios due to prohibitive signaling overhead and coordination complexity. In contrast, in RA (also known as uncoordinated access), users transmit their messages simultaneously as soon as new data is generated, without requiring explicit coordination with the central unit. This approach reduces both latency and signaling overhead. These advantages make RA solutions particularly suitable for next-generation (6G and beyond) communication networks, especially when serving a large number of users [5, 6].
Figure 1: Taxonomy of different MA techniques.
Traditional RA solutions like ALOHA [7] and its variants suffer from excessive collisions as the number of users grows. More sophisticated grant-free RA schemes [8] employ unique signatures or preambles for each potential user to facilitate collision reduction at the physical layer. Grant-free RA solutions can accommodate more users compared to ALOHA, as they can employ user scheduling via a limited amount of feedback [9], or device-specific patterns such as codebook structures and spreading sequences [10]. However, the distinct pilot/preamble requirement for each potential user becomes infeasible as the number of connected devices reaches hundreds of millions. On the other hand, in many of the emerging applications, such as such as massive MTC and IoT, devices become active only sporadically. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer) with low computational and storage capabilities, low duty cycles, and small payloads [3].
To mitigate the challenge of identifying and decoding messages from such a large number of potential devices, Polyanskiy introduced unsourced random access (URA) paradigm [11]. URA fundamentally reimagines the massive access problem by decoupling user identification from data transmission. Instead of assigning unique preambles or codebooks to each potential user employ a common codebook, and the receiver focuses solely on recovering the list of transmitted messages. This paradigm shift drastically reduces system complexity and enables practical implementations for truly massive numbers of devices. Following Polyanskiy’s pioneering work, several low-complexity coding solutions have been developed in the subsequent literature, in which, by employing successive interference cancellation (SIC) techniques, collisions are resolved up to a certain level depending on user transmit powers. Many techniques, such as random spreading [12], multistage orthogonal pilots [13], and time frame slotting [14], have been employed to mitigate the collisions and interference. While a single-antenna receiver in a Gaussian MA channel (GMAC) setup is considered in the earlier works, the literature is shifted to fading channels with multiple-antenna receivers, especially massive multiple-input multiple-output (MIMO), to support more practical channels and an even greater number of users. A block diagram of different MA techniques is depicted in Fig. 1 alongside the three main set-ups in URA research, namely, GMAC, and single-antenna and MIMO fading channels. The figure also shows the way this survey is structured.
Although close to ten years have passed since Polyanskiy’s initial work, and unsourced random access has been studied in many papers, there have been limited efforts to review the existing schemes on URA. The first attempt is a survey on massive access [15], the authors review both many-user MA, where each user is assigned an individual codebook (i.e., the coordinated approach), and URA, which they call uncoordinated massive access. However, since URA was in its early stages at the time of this survey, it mostly covers the earlier works of GMAC URA. In [16], the focus is again on GMAC URA, and the authors review the design paradigms and existing coding schemes for URA over GMAC while fading channels and MIMO URA are presented as open problems. In a less URA-oriented survey [17], an overview of compressive sensing-based grant-free massive access is covered, where some compressed sensing-based URA schemes are also included. In a recent survey [18], through an overview of the existing URA schemes studying mostly GMAC, the authors present URA as a suitable candidate for next-generation communication systems and show that it can support the 3rd Generation Partnership Project (3GPP) standardization.
While the existing surveys focus on the GMAC setup, this survey also covers the fading URA solutions, which comprise the majority of works in the URA literature. More specifically, we conduct a comprehensive review of the existing URA solutions in the literature in its three main setups, namely, URA over GMAC and single-antenna and MIMO fading channels, accompanied by a discussion on the main challenges with possible solutions and many promising future research directions. We study works that put forth coding schemes as well as those that offer fundamental information-theoretical limit calculations for the URA setup. To the best of our knowledge, this is the first comprehensive survey on URA, reviewing this crucial emerging paradigm for 6G and beyond communications in all aspects. A summary of the existing surveys and our contribution can be found in Table I.
Table I: Summary of the surveys on URA.
Paper Scope
Wu et al., 2020 [15] Review of many-user MAC and URA with a focus on GMAC.
Li et al., 2022 [16] Review of basics of URA and coding schemes on URA over GMAC.
Gao et al., 2024 [17] Review of grant-free massive access with a focus on compressed sensing-based schemes.
Liva et al., 2024 [18] Review of URA over GMAC and showing that it can support 3GPP standardization.
This survey A comprehensive review of the URA literature, covering all existing solutions in GMAC and single-antenna and MIMO fading channels.
Table II: List of acronyms
Acronym Definition
5G fifth generation
5G-NR 5G new radio
6G sixth generation
AD activity detection
AMP approximate message passing
AOA angle of arrival
AP access point
AWGN additive white Gaussian noise
BCH Bose–Chaudhuri–Hocquenghem
BP belief propagation
BPSK binary phase shift keying
BS base station
CB-MLD covariance-based MLD
CCS coded/coupled compressive sensing
CDMA code division MA
CP cyclic prefix
CRC cyclic redundancy check
CS compressed sensing
CSA coded slotted ALOHA
CSI channel state information
DFT discrete Fourier transform
DOF degree of freedom
FBL finite blocklength
FDMA frequency division MA
FEC forward error correction
FFT fast Fourier transform
FL federated learning
GMAC Gaussian MA channel
gOMP generalized orthogonal matching pursuit
HWI hardware impairments
i.i.d. independent and identically distributed
IDMA interleave-division MA
IISD iterative inter-symbol decoder
IoT Internet-of-things
IRSA irregular repetition slotted ALOHA
ISAC integrated sensing and communication
LASSO least absolute shrinkage and selection operator
LDPC low-density parity check
LLR log-likelihood ratio
LMMSE linear minimum mean square error
LS least squares
MA multiple access
Acronym Definition
MAC MA channel
MIMO multiple-input multiple-output
ML machine learning
MLD maximum likelihood detection
MMSE minimum mean square error
mMTC massive machine-type communications
MMV-AMP multiple measurement vector AMP
MRC maximal ratio combining
MSE mean square error
MS-MRA multi-stage set-up with multiple receive antennas
MS-MRA-WOPBE MS-MRA without pilot bits encoding
MSUG multi-stage set-up with user grouping
MSUG-MRA MSUG for multiple receive antennas
MTC machine-type communications
MUSIC multiple signal classification
NNLS non-negative least squares
NOMA non-orthogonal MA
NP Neyman-Pearson
ODMA on-off division MA
OFDM orthogonal frequency division multiplexing
OMA orthogonal MA
OMP orthogonal matching pursuit
PMF probability mass function
PUPE per-user probability of error
QPSK quadrature phase shift keying
RA repeat accumulate
RIS reconfigurable intelligent surface
RS Reed-Solomon
SCLD successive cancellation list decoder
SCMA sparse code MA
SDMA space division MA
SIC successive interference cancellation
SINR signal-to-interference-plus-noise ratio
SISO soft-input soft-output
SKP sparse Kronecker product
SNR signal-to-noise ratio
SPARC sparse regression code
TDMA time division MA
TIN treating interference as noise
UE user equipment
UNISAC unsourced ISAC
URA unsourced RA
The remainder of the survey is organized as follows. In Section II, a general discussion of MA techniques is provided. We start with coordinated MA techniques and find our way to random access. In Section III, we focus on URA as a new paradigm for uncoordinated MA communications with a massive number of potential devices, enumerate its advantages, and hint at some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the fading works with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions.
The acronyms used in this paper are alphabetically summarized in Table II.
II Background: Multiple Access
Multiple access refers to the scenario in which two or more users simultaneously share the system resources. In this section, we provide a concise overview of the main MA approaches and enumerate their pros and cons in order to better understand the position of URA solutions in next-generation communication systems.
II-A Coordinated Access
One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combinations) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework are time-division MA (TDMA), frequency-division MA (FDMA), orthogonal frequency-division MA (OFDMA), code-division MA (CDMA), and space-division MA (SDMA). In TDMA and FDMA, non-overlapping time slots and frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission [4]. On the other hand, in CDMA, all the users can simultaneously utilize all the time-frequency resources through different (nearly) orthogonal code sequences coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing [19].
All of the approaches described in the previous paragraph are in the category of OMA, where the number of users is strictly limited due to the orthogonal resource assignment. In order to increase the number of supported users, non-orthogonal MA (NOMA) techniques are investigated in the literature, where multiple users can utilize the same resources to transmit their data, allowing the system to support more users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [20], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through successive interference cancellation (SIC). On the other hand, code-domain NOMA is similar to classical CDMA in the sense that the users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [21]. Here, the receiver can employ message-passing algorithms utilizing the code structure to differentiate the users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division MA [22], low-density spreading aided orthogonal frequency division multiplexing [23], sparse code MA (SCMA) [24], and multiuser shared access [25].
II-B Random Access
All of the schemes described in the previous subsection are coordinated. That is, they require scheduling grants and coordination with the BS before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5-3 according to 3GPP studies, as the system performance significantly degrades beyond that threshold [15]. The common solution addressing these drawbacks is random access, where the users perform their transmission without any coordination by the BS whenever they have data. ALOHA and grant-free random access are the two common RA approaches in the literature.
II-B1 ALOHA & Variants
An early RA approach is the ALOHA protocol [7] (also known as pure ALOHA) proposed in the 1970s, where a device sends the entire packet at a randomly selected time. If no acknowledgment is received within a certain period of time, it is assumed that a collision has occurred with a packet sent by another device, and the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occur or until the process is terminated by the user console. Slotted ALOHA (SA) is a variant of ALOHA, where the packets are transmitted in designated time slots, doubling the maximum throughput compared to the pure ALOHA [26]. However, collision is still a major challenge in SA, as several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 2.
Figure 2: An illustration of pure and slotted ALOHA. Red boxes show the collided packets and green boxes are the successfully transmitted ones.
With the motivation to increase the throughput, several variants of SA are proposed in the recent literature, where the main idea is to transmit multiple copies of a packet in different slots. If it is successfully recovered in one slot, the effect of its replicas in the other slots is removed. Based on this idea, two copies of the same packet is transmitted in contention resolution diversity slotted ALOHA (CRDSA) [27], and the packets are repeated a different number of times rather than twice in irregular repetition slotted ALOHA (IRSA) proposed in [28]. An extension of IRSA exploiting forward error correction that is called coded slotted ALOHA (CSA) [29] approaches a throughput of almost one when sufficiently low coding rates are employed. IRSA and CSA are further studied in [30, 31, 32, 33, 34]. Moreover, the resolution of partial collisions in random access setups is discussed in [35].
II-B2 Grant-Free Random Access
Grant-free RA is another RA candidate to enable communication of a large number of users where the users transmit their data by employing the shared system resources without any prior grant from the BS, which reduces the latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks, which means that there is user identity during the data transmission phase, and the receiver can utilize this information for activity detection or user separation [36]. This is called sourced (grant-free) random access, which one of solutions based on it, grant-free NOMA, is the prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that, the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection, channel estimation, and data recovery, and synchronization with minimal overhead [37].
Among the two main categories of NOMA, the power-domain NOMA is hard to extend to the grant-free scenario as the successful recovery of the transmitted packets at the receiver heavily depends on the power difference of the users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery [38]. For instance, a grant-free version of SCMA is proposed in [39]. Furthermore, the inherent sparsity of the user activity in grant-free MA is utilized in [40, 41, 42, 43] through compressive sensing algorithms for multiuser detection.
III URA: A Recent Paradigm
In some next-generation communications applications, such as massive MTC and massive IoT, the total number of devices can be in the order of millions, and they are only active very sporadically, as illustrated in Fig. 3. This makes achieving any level of coordination infeasible. As a result, none of the MA techniques summarized in the previous section is suitable for enabling the communication of these devices. Specifically, the complexity of the grant-free MA is prohibitive due to the codebook size when each potential user is assigned a unique preamble. On the other hand, ALOHA leads to excessive collisions, significantly degrading communication efficiency.
To address this problem, Polyanskiy proposed the URA paradigm in [11] where the data transmission and user identification problems are decoupled via the utilization of a common codebook by all users. So, the receiver is not interested in user identities during the data transmission, with the decoding being done only up to a permutation of the transmitted messages.
III-A General System Model
The URA system model in a synchronous setup as follows. Considering a URA system with an $M_{r}$-antenna receiver and single-antenna users, in which $K_{a}$ out of $K_{T}$ users are active at a given time frame of length $n$, the received signal, $\mathbf{Y}\in\mathbb{C}^{M_{r}\times n}$, in the absence of synchronization errors can be written in the following general form [13]
$\mathbf{Y}=\sum_{i=1}^{K_{a}}\mathbf{h}_{i}{\mathbf{x}_{i}(\mathbf{u}_{i})+{\mathbf{Z}}}\in\mathbb{C}^{M_{r}\times n},$ (1)
with the following common models in the URA literature
$\begin{cases}M_{r}=1,\mathbf{h}_{i}=1,&\text{GMAC}\\ M_{r}=1,\mathbf{h}_{i}\sim\mathcal{CN}(0,1),&\text{single-antenna fading channel}\\ M_{r}>1,\mathbf{h}_{i}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{r}}),&\text{MIMO fading channel}\\ \end{cases}$ (2)
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the length-$n$ transmitted signal of the $i$-th user corresponding to its message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of length $B$, and ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ is the circularly-symmetric additive white Gaussian noise (AWGN) whose elements are drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$, with $\sigma_{z}^{2}$ being the power of noise. In the literature, it is assumed that each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine a list of transmitted messages, $\mathbf{u}_{i}^{\prime}s,\ \forall i=1,2,...,K_{a}$, by processing the received signal $\mathbf{Y}$.
Figure 3: An illustration of a system where a large number of potential devices (users) try to communicate with an access point (AP), with only a small subset of them being active at each time frame.
In the URA literature, the common setup involves users having messages of fixed length $B$ bits from a set of the possible messages $[M]$ with cardinality $M$, and the objective is to minimize the required energy-per-bit for a target per-user probability of error (PUPE), as the performance metric. The energy-per-bit can be written as
$\dfrac{E_{b}}{N_{0}}=\dfrac{nP}{c\sigma_{z}^{2}B},$ (3)
where $P$ denotes the average power of each user per channel use, i.e., $\norm{\mathbf{x}_{i}}^{2}\leq nP$, and we have $c=2$ for real-valued and $c=1$ for complex channels. Also, PUPE is defined in terms of the probability of false-alarm, $p_{fa}$, and the probability of misdetection, $p_{md}$, as
$P_{e}=p_{fa}+p_{md},$ (4)
where $p_{md}=\dfrac{1}{K_{a}}\sum_{i\in\mathcal{K}_{a}}{\mathrm{Pr}(\mathbf{u}_{i}\notin\mathcal{L}_{d})}$, where $\mathcal{K}_{a}$ and $\mathcal{L}_{d}$ are the set of active user indices and the list of decoded messages, respectively, and $p_{fa}=\mathbb{E}\left\{\dfrac{n_{fa}}{|\mathcal{L}_{d}|}\right\}$, with $n_{fa}$ being the number of decoded messages that were indeed not sent.
III-B Challenges
In the following sections, we review the existing URA solutions in the literature. However, in order to have a better understanding of the URA setup and the reasoning behind the provided solutions in the literature, we first need to know the main challenges faced in a URA setup, which will be hinted at in the rest of this section.
Interference: Interference stands out as the primary challenge in MA channels, especially random access due to less coordination. It has an even more detrimental effect on the URA setup since there is no coordination between the receiver and the users. Failure to incorporate strategies for reducing or canceling it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we analyze the treating interference as noise (TIN) scheme, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. The signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy for a GMAC scenario can be written as
$\alpha_{i}=\dfrac{P}{\sigma_{z}^{2}+{(K_{a}-1)}P},$ (5)
[…]
$\alpha_{i}\approx\dfrac{1}{K_{a}}.$ (6)
This equation illustrates that when the number of users surpasses a specific threshold (more specifically, when it is much greater than the inverse of transmit signal-to-noise ratio (SNR), i.e., $K\gg\frac{\sigma^{2}_{z}}{P}$, which is not a very large value in many envisioned URA applications), the SINR becomes inversely proportional with $K_{a}$. So, boosting the transmit power does not result in an enhanced SINR or, equivalently, the system performance. Consequently, in systems such as URA, the relatively large number of active users significantly rise in $K_{a}$ (equivalent to an increase in interference) adversely impacts the system performance.
There are two main techniques in the URA literature to alleviate the effects of interference, with the most common one being SIC. SIC is a useful technique for interference mitigation in wireless multiuser scenarios, which removes the effects of successfully detected messages from the received signal to decrease the interference before another round of decoding. Especially in URA, SIC is an essential step in achieving low error rates due to heavy multiuser interference. The other technique for interference management involves randomly separating users into different groups, e.g., separation in time by slotting the time frame, in the power domain by employing different transmit power levels, or in the code domain by multiplying or concatenating a known sequence to the data segment.
  1. 1.
    Time-domain separation: Dividing the time frame into multiple slots and letting each user randomly select a slot to transmit its data can lead to enhanced system performance. In a slotted structure, the trade-off is between the effects of decreased interference due to the reduced number of interfering users in each slot and decreased channel coding power due to the fewer channel uses available to each user. Therefore, optimizing the number of slots as a function of system parameters such as active user load plays an important role in benefiting from the slotting.
  2. 2.
    Code-domain separation: In some URA schemes, each user randomly selects a preamble/pilot/pattern and uses it to prepare its data for transmission. One way is to transmit a concatenation of a preamble or pilot and the main signal. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. Then, minimum mean square error (LMMSE) or maximum ratio combining (MRC) solutions can be employed to separate the user packets, before decoding them using single-user channel decoders. Another way is to multiply a randomly selected pattern with the main signal, which is used in random-spreading-based solutions. Due to limited available power and time resources, employing these techniques usually decreases the amount of power and time resources dedicated to the data (main signal) part. So, in order for them to be effective in interference management, the choices of type, length, and power of the pilots and pattern sequences are of great importance.
  3. 3.
    Power-domain separation: An alternative approach for user separation, especially effective in GMAC scenarios, is the so-called power diversity technique [44], where each user randomly picks a transmit power level from a limited set with a certain probability distribution, satisfying the overall power constraint. Since the users are now separated by different power levels, the receiver, with a high probability, decodes the users with the highest power level first, and after each SIC, it moves to decode the lower power levels. The appropriate design of the power diversity parameters, namely, the number of power levels, their values, and their probability distribution, is very important in achieving the power diversity gain.
Collisions: In many classic random access setups, such as ALOHA, collision is used interchangeably with interference, i.e., when user packs are fully or partially received at the same time at the receiver. However, in the URA literature, collision by default means the collision of either user messages or user pilots, which are when two or more users have exactly the same messages or pilot sequences, respectively.
  1. 1.
    Message collision: In the case of user message collision, with the help of a randomized encoding process independent of user messages, the receiver can recover collided user messages to some extent. Note that in a conventional URA setup, since the receiver is only interested in the list of transmitted messages, being able to recover at least one of the collided messages does not increase the misdetection rate. However, this matter becomes more important in some URA systems where the receiver is also interested in the frequency of user messages.
  2. 2.
    Pilot collision: Many existing fading URA solutions employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users independently pick a preamble/pilot from a common codebook, either randomly or based on a part of their message bits. Hence, if the pilot bits of multiple users are the same (called pilot contamination in massive MIMO literature [45]), a collision occurs, and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the pilot collision probability lower than a desired threshold or develop techniques for its mitigation.
Computational Complexity: In order to increase the number of active users supported by a URA system, larger values of parameters such as time frame length and the number of antenna elements (in MIMO systems) are necessary. However, the computational complexity of URA solutions escalates superlinearly with these parameters as well as the number of active users Consequently, solutions for a URA system with a high dimension, temporal (time frame length), spatial (number of receiver antennas), and/or the number of active users, are anticipated to exhibit a high computational complexity. To mitigate this complexity, various strategies are employed in the literature, such as slotted transmissions and coded compressed sensing (CCS). The idea of CCS is to divide the bit sequence into smaller segments, map each segment to a column of a common codebook, and stitch them using a tree code instead of mapping the entire bit sequence to the columns of a $2^{B}\times n$ codebook, which would lead to a prohibitive receiver complexity. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In this section, we summarized the main challenges in a URA setup along with a high-level discussion on the proposed solutions in the literature. In the two subsequent sections, we take a more detailed look at these solutions, with Section IV dedicated to GMAC and Section V to single-antenna (V.A) and MIMO (V.B) fading setups. We take a systematic approach in presenting the URA solutions in these sections: different contributions are grouped based on their solution approach. In each group, the underlying approach is explained with an emphasis on the main work accompanied by a block diagram of its encoder and decoder. The studies are presented in chronological order. To provide an overall view, at the end of each main subsection, a summary of the primary contributions is presented, and a comparative discussion of their performance is provided, accompanied by some numerical results.
IV URA over GMAC
In this section, we present an overview of GMAC URA. To this end, we first present an achievability bound on the performance of the GMAC URA, as presented in Polyanskiy’s original paper [11], and then, review the existing GMAC URA solutions to achieve it.
IV-A Performance Bounds
In [11], Polyanskiy introduces the URA paradigm and obtains an achievability bound on its performance on the following GMAC model with single-antenna users and receiver:
$\mathbf{y}=\mathbf{x}_{1}+\dots+\mathbf{x}_{K_{a}}+\mathbf{z},$ (7)
where $K_{a}$ is the number of active users (i.e., the active user load), $\mathbf{x}_{i}$ is the transmitted signal of the $i$-th user, $\mathbf{y}$ is the received signal and $\mathbf{z}$ is the circularly-symmetric AWGN noise vector with independent and identically distributed (i.i.d.) elements, i.e., $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$.
In [11], an $(M,n,\epsilon)$ RA code, where $\epsilon$ is the target PUPE, is defined as follows
Definition 1.
An $(M,n,\epsilon)$ RA code for the $K_{a}$-user channel $P_{Y\mid X^{K_{a}}}$ is a pair of (possibly randomized) maps, the encoder $f:[M]\rightarrow\mathcal{X}^{n}$ and the decoder $g:\mathcal{Y}^{n}\rightarrow\binom{[M]}{K_{a}}$, satisfying
$\frac{1}{K_{a}}\sum\limits_{j=1}^{K_{a}}\mathbb{P}[E_{j}]\leq\epsilon,$ (8)
where $E_{j}\triangleq\{W_{j}\notin\mathcal{L}\}\cup\{W_{j}=W_{i}\hskip 5.69054pt\text{for some }i\neq j\}$ is the error event for the $j$-th user, $W_{1},\dots,W_{K_{a}}$ are independent and uniform on $[M]$ messages, and $\mathcal{L}$ is the list of the estimated user messages that is produced by the decoder. Note that this error definition consists of the misdetection error and the error resulting from users having the same message sequence. Firstly, it does not take into account the false alarm probability. The reason is that since the number of active users is assumed to be known and a fixed-size decoding list of length $K_{a}$ is used at the receiver in [11], the false alarm probability becomes the same as the misdetection error, so there is no need for its inclusion in the error definition. However, the number of active users is not known by the receiver in practice, and needs to be estimated as is done in other URA solutions, either explicitly or implicitly. Therefore, for a more complete analysis of URA systems, both misdetection and false alarm errors need to be taken into account in the error definition. Another point is that it is assumed in this error definition that any instance of two users having the same message leads to an error, which is not valid. However, this assumption does not violate the achievability of the proposed bound as it assumes a higher error.
The following achievability bound has been used as a benchmark for the energy efficiency of URA over a real-valued AWGN channel.
Theorem 1 (Random coding achievability bound [11]).
Fix $P^{\prime}<P$. There exists an $(M,n,\epsilon)$ random-access code for $K_{a}$-user GMAC satisfying power-constraint $P$ and
$\epsilon\leq\sum\limits_{i=1}^{K_{a}}\frac{t}{K_{a}}min(p_{t},q_{t})+p_{0},$ (9)
where $p_{0}=\frac{\binom{Ka}{2}}{M}+K_{a}\mathbb{P}\Bigg{[}\frac{1}{n}\sum\limits_{j=1}^{n}Z_{j}^{2}>\frac{P}{P^{\prime}}\Bigg{]}$, $p_{t}=e^{-nE(t)}$, $E(t)=\max\limits_{0\leq\rho,\rho_{1}\leq 1}-\rho\rho_{1}tR_{1}-\rho_{1}R_{2}+E_{0}(\rho,\rho_{1})$, $E_{0}=\rho_{1}a+\frac{1}{2}\log(1-2b\rho_{1})$, $a=\frac{\rho}{2}\log(1+2P^{\prime}t\lambda)+\frac{1}{2}\log(1+2P^{\prime}t\mu)$, $b=\rho\lambda-\frac{\mu}{1+2P^{\prime}t\mu}$, $\mu=\frac{\rho\lambda}{1+2P^{\prime}t\lambda}$, $\lambda=\frac{P^{\prime}t-1+\sqrt{D}}{4(1+\rho_{1}\rho)P^{\prime}t}$, $D=(P^{\prime}t-1)^{2}+4P^{\prime}t\frac{1+\rho\rho_{1}}{1+\rho}$, $R_{1}=\frac{1}{n}\log M-\frac{1}{nt}\log(t!)$, $R_{2}=\frac{1}{n}\log\binom{K_{a}}{t}$, and $q_{t}$ is defined as
$q_{t}=\inf_{\gamma}\mathbb{P}[I_{t}\leq\gamma]+exp(n(tR_{1}+R_{2})-\gamma),$ (10)
where the random variable $I_{t}$ is defined as
$I_{t}=\min_{S_{0}}i_{t}(c(S_{0});Y|c(S_{0}^{c}),$ (11)
where $c(S_{0})\triangleq\sum\limits_{j\in S_{0}}c_{j}$ for $c_{1},\dots,c_{M}\sim\mathcal{N}(0,P^{\prime})$ and the minimum is over all $t$- subsets of $[K_{a}]$. The information density $i_{t}$ is defined as
$i_{t}(a;y|b)=nC_{t}+\frac{\log e}{2}\Big{(}\frac{\norm{y-b}_{2}^{2}}{1+P^{\prime}t}-\norm{y-a-b}_{2}^{2}\Big{)},$ (12)
where $C_{t}=\frac{1}{2}\log(1+P^{\prime}t)$.
The idea behind the proof of this theorem is as follows: Let $S=\{W_{1},\dots,W_{K_{a}}\}$ be a random $K_{a}$-subset of $[M]$. We can bound the error event $F_{t}\triangleq\{\lvert S\setminus\hat{S}\rvert=t\}$ with two different methods. The first one is a Gallager-type bound, defining the error events using the generic subsets of size $t$ in $[K_{a}]$ and $[M]\backslash[K_{a}]$, respectively, using the Chernoff bound and applying Gallager’s $\rho$-trick twice. The second one utilizes the information density in (12), considering the event $\tilde{F}=\{I_{t}<\gamma\}$ with the fixed $\gamma$ and $I_{t}$ is given in (11), and taking union over $S_{0}$.
The random coding bound is an information-theoretic performance limit, and it is derived without any complexity constraints as it assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound. In order to illustrate the ineffectiveness of common RA solutions in a URA setup, we first compare their performance with Polyanskiy’s URA achievability result.
In slotted ALOHA, a user packet is successfully recovered if there is no collision and the single-user decoding is successful, as the collided packets are discarded. Moreover, using the normal approximation in [46], the finite blocklength capacity of TIN can be approximated as
$\log M\approx nC_{TIN}(P)-\sqrt{\frac{nP\log^{2}e}{1+K_{a}P}}Q^{-1}(\epsilon),$ (13)
where $C_{TIN}(P)=\frac{1}{2}\log\left(1+\frac{P}{1+(K_{a}-1)P}\right)$, $Q^{-1}(\cdot)$ denotes the inverse Q-function.
In [11], the performances of these approaches are evaluated assuming that there are $K_{a}$ active users transmitting $B=100$ bits of information through a transmission time frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by subsequent GMAC URA works for performance comparison, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
The required $E_{b}/N_{0}$ versus number of active users for these schemes are plotted in Fig. 4 along with the random coding bound derived in [11] and a converse bound for $K_{a}$-GMAC from [47]. The results in Fig. 4 show that the conventional RA techniques (i.e., slotted ALOHA and TIN) are energy inefficient in a URA setup as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. The performance of the grant-based TDMA scheme is also plotted in the figure (referred to as Orthogonal M.A) as a benchmark, in which the time frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame, with its required $E_{b}/N_{0}$ being calculated using the normal approximation in [46]. However, note that the TDMA scheme is not feasible in a URA setup, mainly due to the lack of coordination. Putting the coordination issue aside, the system works in a regime of $n<K_{a}B$ in some URA applications, which renders orthogonal solutions infeasible.
Figure 4: Performance comparison of conventional random access techniques with Polyanskiy’s URA (random coding) achievability bound.
The random coding bound in [11] is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors of [48] extend this bound to the case that the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [49]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [50], where the alarm message needs to be decoded with a lower error rate than the standard ones. This result is extended to the case of asynchronous GMAC in [51], where the delay in transmission is assumed to be up to a small portion of the time frame. However, the provided achievability bound is a loose bound for the asynchronous case as the authors have employed a fixed time frame decoder in which all the packets that have some overlaps with the end of the time frame are deemed undecodable, leading to huge performance degradation, which is not valid in a fully asynchronous setup, where the time horizon is unbounded. Furthermore, the problem of joint delay and activity detection in the asynchronous setup is investigated in [52] using methods of tropical geometry, which is the study of the geometric properties of polynomials when addition is replaced by minimization and multiplication by ordinary addition. Moreover, the conditions for path delays for zero-error detection are identified.
IV-B URA Solutions for GMAC
As shown in the previous subsection, the conventional RA techniques are energy-inefficient for the URA setup. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems. The main approaches for this purpose in the literature can be categorized as slotting the transmission frame, coded compressed sensing (CCS), random spreading, sparse spreading, and on-off division MA (ODMA). In the following, we review the existing practical coding schemes for URA over GMAC based on this classification.
IV-B1 Slotting-based schemes
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [53], where a concatenated coding scheme is considered combined with the $T$-fold ALOHA protocol. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in the slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [53], each user randomly chooses one slot to transmit its message. This concatenated coding scheme consists of an inner binary linear code which is used to decode the modulo-2 sum of the codewords transmitted within a slot and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [54], where the authors employ a concatenated coding scheme with polar and BCH codes as the inner and outer codes, respectively, and propose a decoding approach exchanging the information between inner and outer decoding.
Figure 5: The slotted transmission structure in [55] based on $T$-fold IRSA.
The authors of [55] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as T-fold IRSA. They propose to split the message into two parts and employ a coding scheme based on a combination of compressed sensing (CS) [56] and LDPC coding at the slot level. Namely, the users pick a preamble determined by the first part of the message from a common codebook, and the second part is encoded by an LDPC code and interleaved before the transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by the non-negative least squares (NNLS) algorithm and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. The general transmission structure in T-fold IRSA based URA solutions is depicted in Fig. 5.
$T$-fold IRSA is also utilized in [57] by replacing the LDPC codes with the polar codes and utilizing joint successive cancellation decoding, which provides a significant performance gain over [55]. In [58], the authors employ random signatures to convert the equal-gain channel into a more favorable i.i.d. fading scenario along with trellis-based codes in a $T$-fold ALOHA structure. Their proposed approach provides a low-complexity solution with moderate performance. Moreover, the authors of [59] employ finite-field coding in a slotted structure. More specifically, the users encode their data using a non-binary LDPC code and transmit it after repetition, zero-padding, and interleaving. At the receiver, intra-slot decoding is performed by a joint BP algorithm over a non-binary factor graph. Their proposed approach improves the performance of the slotted schemes in high user loads, compared to its previous counterparts.
IV-B2 CCS-based schemes
Figure 6: Encoding-decoding structure in a coded compressed sensing (CCS) approach.
Figure 7: A typical decoding structure in a random spreading approach.
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [60, 61, 62, 63, 64, 65, 66], where the messages are divided into sub-blocks and each sub-block is first encoded by an outer code by appending redundant parity bits before being mapped to a CS codebook for the inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer tree code. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 6. More specifically, the encoding process is performed in two phases: 1) In the tree encoder, $B$ bits of information are divided into $S$ parts, and all of these segments, except the initial one, are appended by additional parity-check bits; 2) in the CS encoder, each of $S$ sub-message is mapped to a CS codebook to select a sequence for transmission. Hence, the received signal is comprised of $S$ slots, with each slot being a superposition of the sequences from $K_{a}$ users. The decoder also functions in two stages: 1) In the initial stage, $K_{a}$ sequences associated with the $K_{a}$ active users are identified within each slot. 2) in the second stage, called outer decoder or tree decoder, with the assistance of the unique parity bits, the sub-messages corresponding to the detected sequences are combined to create a message of length $B$.
Figure 8: Encoding process of sparse IDMA [72].
Several variations of the CCS approach are examined in the URA literature. Specifically, in [60], the sub-blocks are first encoded by a binary linear code and then mapped to the columns of a common sensing matrix for inner encoding. In [61], sparse regression codes (SPARCs) are exploited as inner codes with approximate message passing (AMP) decoding at the receiver. The authors of [62] improve the performance of the proposed scheme in [61] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [63], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is employed for inner decoding, with the addition of tree decoding. Furthermore, the authors of [64] adapt constant-weight codes as the outer code with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is employed as the outer code in [65]. Finally, in a recent work [66], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the remaining solutions for $K_{a}\geq 225$.
IV-B3 Random spreading-based schemes
In another design strategy for URA over GMAC, there is no slotting and the active users utilize the whole transmission frame. In some works following this strategy, the idea is to spread the channel-coded bits through the transmission time frame using random signatures. The random spreading idea was first introduced in [67], where the message is divided into two parts: the first part determines the signature sequence while the second part is encoded by a polar code. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading-based solutions is illustrated in Fig. 7.
In [68], the authors modify the scheme in [67] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. Moreover, [44] improves the performance of the scheme in [67] by letting the users to randomly select a power level from a pre-defined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher power levels are decoded first with a high probability, and the ones with lower power levels are decoded next after SIC. This results in a performance improvement over [67] for $K_{a}\geq 150$. The authors of [69] replace the polar codes in [67] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach improves the decoding performance since the information about the partially decoded codewords can be utilized, and it provides the best performance among the schemes based on random spreading. Furthermore, enhanced spread spectrum ALOHA (E-SSA) [70], an asynchronous RA protocol combining ALOHA with direct sequence spread spectrum is adapted for URA over GMAC in [71]. The authors employ a wrap-around version of E-SSA in order to adapt it for a synchronous URA setup.
IV-B4 IDMA-based schemes
In random spreading based solutions, each active user occupies the entire time frame by random spreading of channel-coded bits for data transmission. In some other works, the active users distribute their codeword bits onto the transmission time frame in different ways. For instance, in the sparse interleave-division MA (IDMA) approach [72], the message bits and the time frame are divided into parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The rest of the bits are encoded by an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted along the rest of the channel frame. This encoding process is illustrated in Fig. 8. At the receiver, after decoding the preamble part via the least absolute shrinkage and selection operator (LASSO) technique, a single joint Tanner graph is utilized for message recovery, exploiting the information of the location of transmitted symbols from the identified preamble.
A similar approach to the sparse IDMA in [72] is adopted in [73] employing polar codes, however, different code lengths are used among the users and TIN is utilized as the decoding algorithm. Moreover, in [74], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts, with the first part being mapped to a column of a common sensing matrix, which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part followed by a multiuser decoding algorithm for the second part. Furthermore, the authors of [75] propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components.
Table III: Main URA solutions for GMAC
Paper Approach Contribution and Techniques
Polyanskiy, 2017 [11] - information-theoretic formulation of URA, and Random coding achievability and converse bounds
Ordentlich et al., 2017 [53] slotting concatenated coding combined with $T$-fold ALOHA
Vem et al., 2019 [55] slotting LDPC coding, preamble transmission, and SIC
Marshakov et al., 2019 [57] slotting preamble transmission, polar coding, and SIC
Amalladinne et al., 2020 [60] CCS inner CS and outer tree coding
Amalladinne et al., 2022 [62] CCS interaction between the inner AMP and the outer BP
Nassaji et al., 2024 [66] CCS eliminating the outer tree code
Pradhan et al., 2022 [72] sparse spreading preamble transmission, LDPC coding and SIC
Han et al., 2021 [75] sparse spreading convolutional coding and SIC
Pradhan et al., 2020 [67] random spreading polar coding and SIC
Ahmadi et al., 2020 [44] random spreading power diversity, polar coding, and SIC
Pradhan et al., 2021 [69] random spreading LDPC coding, soft cancellation, and SIC
Yan et al., 2023 [76] ODMA repeat-accumulate coding and soft cancellation
Ozates et al., 2024 [78] ODMA polar coding and SIC
Yan et al., 2024 [79] ODMA non-binary LDPC coding and soft cancellation
IV-B5 ODMA-based schemes
Another way to distribute the user codeword bits onto the transmission time frame is proposed in [76], where the authors employ on-off division MA (ODMA) [77] in the context of URA. In this approach, the active users exploit a small disjoint portion of the transmission time frame to transmit their data encoded by a repeat-accumulate code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors of [78] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [79], which improves the scheme in [76] by employing a non-binary LDPC code.
Figure 9: Performance comparison of the coding schemes in URA over GMAC for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature.
IV-C Summary and Performance Comparison
A summary of main URA solutions for GMAC can be found in Table III, accompanied by a comparison of their performance in terms of energy efficiency in Fig. 9 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. The results in Fig. 9 demonstrate that random spreading-based schemes in [67, 44, 69] offer good performance for a lower number of active users; however, their performance degrades with the increasing number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA-based approaches in [76, 78, 79] outperform the others except the solution proposed in [66] in the high multiuser interference regime due to its sparsity, while it performs well in the low multiuser regime when combined with polar coding. Moreover, CCS-based approaches can exhibit very good performance for a high number of active users when the rate loss due to the redundant parity bits is alleviated, as shown in [66]. Compared to the achievability bound, the proposed schemes can perform similar or slightly better for $K_{a}\leq 100$, and the gap is a maximum of 0.8 dB for $K_{a}\leq 300$.
IV-D Computational Complexity
In terms of complexity, slotting is the most beneficial approach among the existing solutions for GMAC URA, as the complexity of the operations depending on the user count super-linearly increases by the number of active users in the slot rather than $K_{a}$. In addition, slotting leads to smaller preamble/pilot length and channel code length $n_{c}$, decreasing the complexity of the operations depending on these parameters. CCS also has a low complexity since slot-wise compressed sensing and a low-complexity tree code are employed in conjunction. Namely, its complexity order is linear or quadratic in $K_{a}$ depending on the regime [60]. ODMA is another low-complexity approach, as only a small fraction of the time frame is occupied by each user, which simplifies the iterative multiuser decoding process [76]. More specifically, its complexity order linearly depends on $n_{c}$ and $K_{a}$. On the other hand, random spreading-based schemes where each user occupies the entire time frame have the highest complexity, as the complexity order is cubic in $K_{a}$ due to the MMSE filtering [67]. On another note, the number of computations needed for some decoding operations depends on the transmission time frame length rather than the slot or code length as in slotting or ODMA, which increases the complexity.
V URA over Fading Channels
Since GMAC is not a very realistic channel model for wireless communications, most works in the URA literature consider fading channels, the majority of them dedicated to Rayleigh fading channels. In this section, we first review the bounds on the performance of fading URA solutions that have been provided in the literature. Then, we cover the URA solutions that present low-complexity coding schemes to achieve those bounds. To this end, in order to explain the fading URA approaches in a simpler setup for better understanding, we first cover the single-antenna fading solutions, then, we turn our attention to the MIMO fading ones.
V-A Performance Bounds
In [85, 83, 84, 80], the authors derive some performance bounds for the URA setup over a Rayleigh fading channel. In one of the main analytical results, the authors of [80] present a converse bound for fading URA based on the results in [81] and the meta-converse bound in [82], as stated below.
Theorem 2 (Converse bound [80]).
Let
$L_{n}=n\hskip 2.84526pt\log(1+PG)+\sum\limits_{i=1}^{n}\left(1-\absolutevalue{\sqrt{PG}Z_{i}-\sqrt{1+PG}}^{2}\right),$ (14)
$S_{n}=n\hskip 2.84526pt\log(1+PG)\sum\limits_{i=1}^{n}\left(1-\frac{\absolutevalue{\sqrt{PG}Z_{i}-1}^{2}}{1+PG}\right),$ (15)
where $G=\norm{h}^{2}$ and $Z_{i}\sim\mathcal{CN}(0,1)$. Then, for every $n$ and $10<\epsilon<$ 1, any $(M,n-1,\epsilon)$ code for the quasi-static $K_{a}$ MA satisfies
$\log(M)\leq\log(K_{a})+\log\frac{1}{\mathbb{P}(L_{n}\geq n\gamma_{n})},$ (16)
where $\gamma_{n}$ is the solution of $\mathbb{P}\left[S_{n}\leq n\gamma_{n}\right]=\epsilon$.
For the proof, first, consider the fact that the converse bound when channel state information (CSI) is available at the receiver is a converse for the case with no CSI case as well. Moreover, by symmetry, it is enough to lower bound the probability that a particular user’s message is not in the decoded list. Finally, assuming that the decoder knows the codewords of all the other users (which makes this in essence a single-user bound), the rest of the proof becomes quite straightforward, which can be found in [80]. The authors numerically show that the $E_{b}/N_{0}$ requirements for a fading channel are necessarily higher compared to the AWGN channel. They also mention that, for the common system parameters in a fading URA setup, a Fano-type converse bound is worse than the one presented in Theorem 2.
In [83] and [84], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. The authors in [80] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [85], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering slotted structure and perfect SIC in their analysis, which is tighter than the other existing bounds. However, since the channel coefficients used in a SIC process are estimated, hence imperfect, the assumption of perfect SIC is not realistic, which makes the result in [85] an approximation result rather than an achievability one.
Moreover, in [86, 87], achievability and converse bounds are obtained for URA over MIMO Rayleigh fading channels, whose ML-based achievability bound is obtained by employing a projection decoder presented in [88].
V-B Single-Antenna Fading Channels
We first take a look at the existing URA fading solutions that assume a single-antenna receiver. The works in [72, 94, 96, 85, 80, 89, 90, 92, 95, 93, 91] introduce practical coding schemes tailored for the URA scheme on single-antenna fading channels. Within this collection of works, some assume that the received signals from all users are synchronized [72, 85, 80, 89], while others take into account asynchronous transmission [95, 90, 92, 93, 91]. To estimate the delay of asynchronous signals, [95, 90, 92, 91] use OFDM transmission to convert the time delay to the phase shift in the frequency domain, while the authors in [93] consider time-domain transmission where the delay of a signal is detected via correlation-based approach.
Focusing on the Rayleigh fading channels, in the following, we present the existing fading URA solutions, categorized based on their general approach.
Figure 10: Another representation of encoding (top) and decoding (bottom) structures of a CCS scheme with 4 slots [60].
V-B1 CCS-based schemes
As explained in the previous section, CCS was originally proposed for URA over GMAC. Nevertheless, it has been applied to the fading scenario as well in several works [89, 93, 96]. In Fig. 10, another representation of the encoding (upper half of the figure) and decoding (lower half) structures of a CCS-based scheme with four slots are depicted, where each user’s signal is divided into four parts, with each one being appended to a parity-check bit, which is then mapped to a codebook to construct the pilot of the user in a slot. Also, at the receiver, sub-messages of different users are detected at each slot, and corresponding ones in different slots are stitched together to form the complete message sequence of a user. In [89], the authors modify the tree decoder of the GMAC CCS scheme in [60] by replacing it with two codes capable of correcting $t$ errors, in which the first one modifies the conventional tree code by giving it the ability to recover $t$ errors (called $t$-tree), and the second one is a Reed-Solomon (RS) code. In [96], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS encoding, with AMP at the receiver. Moreover, in [93], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all potential delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays when performing the CS decoding.
V-B2 Joint decoding-based schemes
In [80], a joint decoding approach is employed. For encoding, this scheme partitions the time frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers two significant benefits: it reduces the computational load on the decoder by reducing the length of the LDPC signal, and mitigates per-slot interference by distributing users across slots. This scheme utilizes an iterative belief propagation (BP) decoder, as illustrated in Fig. 11. The graph incorporates four distinct node types: variable nodes (in red), and check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to simultaneously update the estimated values of both the channel coefficients and the symbols for each user.
Figure 11: Iterative BP decoder for a two-users LDPC-based scheme with joint decoding [80].
V-B3 TIN-SIC-based schemes
A lower-complexity alternative to joint decoding is TIN-based decoding. Similar to GMAC, the TIN solution for fading channels operates under the assumption of the presence of just a single user and treats other interfering users as noise. Consequently, this approach results in an exceptionally low average SNR due to the fact that it encounters interference-plus-noise as the system’s overall noise level, leading to a very low performance. However, when combined with an appropriate channel code and SIC, it can lead to good performance.
Table IV: A summary of URA solutions for fading channels with a single-antenna receiver.
Paper Approach Description Synchron. vs. Asynch.
Kowshik et al., 2019&2021 [83, 84] - Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron.
Andreev et al., 2019 [90, 91] TIN-SIC A new coding scheme containing LDPC code, OFDM, slotting, and SIC. Asynch.
Andreev et al., 2020 [85] TIN-SIC Non-asymptotic achievability and converse bounds for unknown CSI + a new coding scheme containing slotting, polar code, and SIC. Synchron.
Kowshik et al., 2019&2020 [80] joint decoding Achievability and converse bounds for known and unknown CSI + a scheme containing LDPC code, slotting, and iterative joint decoding. Synchron.
Amalladinne et al., 2019 [93] CCS-based They extend CCS in [60] to accommodate asynchronous scenarios. Asynch.
Andreev et al., 2022 [89] CCS-based Improving CCS in [60] by replacing the conventional outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron.
Yun et al., 2024 [96] CCS-based A slotted scheme including CCS and AMP in a block fading model. Synchron.
Nassaji et al., 2022 [94] Spreading-based A non-slotted scheme including, preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron.
Chen et al., 2017 [92] OFDM+ SIC A new coding scheme containing OFDM and SIC. Asynch.
Ozates et al., 2023 [95] OFDM+SIC A new coding scheme in the frequency-selective channel, containing OFDM, pilot transmission, slotting, polar code, and SIC. Asynch.
TIN-SIC with LDPC coding: In [90], building upon their asynchronous schemes presented in [80], the authors present a low-complexity coding scheme which incorporates orthogonal frequency division multiplexing (OFDM) techniques to mitigate the effects of the asynchronous transmission by converting the time delays into phase shifts, and employing SIC to decode the user packets. More numerical results for this scheme is presented in [91].
TIN-SIC with Polar Coding: In [85], a TIN-SIC solution in conjunction with polar coding is presented, where the time frame in this approach is divided into $S$ slots. Each user randomly selects a slot to transmit its polar coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by an SIC. This is because different users in a slot are prone to different fading coefficients; some of them with larger coefficients are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating others as noise, its channel coefficient is estimated using the orthogonal matching pursuit (OMP) technique. Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is demonstrated in Fig. 12.
Figure 12: Encoding and decoding processes of TIN-SIC scheme with polar codes in [85] in a slot with $K$ users.
V-B4 Alternative solutions
There are also some other solutions focusing on URA over fading channels with a single-antenna receiver. For instance, in [94], the active users utilize the entire transmission time frame to transmit their data. Specifically, the time frame is divided into pilot and data parts. The active users pick a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits along the rest of the transmission time frame with permutation and scrambling in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits. Furthermore, the authors of [95] study URA over frequency-selective channels and propose OFDM to mitigate adverse channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
Figure 13: Performance comparison of the coding schemes in fading URA with a single-antenna receiver, with a target PUPE of $0.1$, $n=30000$, and $B=100$ (the common set of parameters in the literature).
Summary and Performance Comparison:
A summary of URA studies over fading channels involving a single-antenna receiver is given in Table IV. Moreover, the performance of different coding schemes for the single-antenna fading URA setup is compared in Fig. 13 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The reason that some diagrams seem not to be fully plotted is that they diverge after a certain active user load, i.e., they cannot achieve the target PUPE no matter how much the $E_{b}/N_{0}$ is increased. As can be observed from Fig. 13, the CCS-based scheme in [89] exhibits lower energy efficiency compared to the other ones since it employs a sub-optimal coding scheme, while others use strong channel codes such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in [80] is inferior to that of the ones in [72], [91] and [85], which can be attributed to the absence of an efficient SIC block in this LDPC-based scheme. On the other hand, the TIN-SIC approach in [91] and [85] offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when it is combined with polar coding as in [85]. Nevertheless, the TIN-SIC scheme in [85] is outperformed by the spreading-based scheme in [94], and the ODMA-based scheme employing polar codes in [78], which offers the best performance among the proposed schemes, due to the sparsity of the transmissions.
Computational Complexity:
We now compare the computational complexity of the single-antenna fading URA solutions described in this subsection. The tree-code-based CCS approach has a low complexity, which is linear in the list of messages at each decoding step and exponential in the number of message bits in each slot [89]. In addition, the decoding operations in [89] can be implemented by additions, which are easier to implement in hardware compared to multiplications. The complexity of the $T$-fold ALOHA-based schemes employing LDPC with joint decoding [80] or polar coding with TIN-SIC [85] is mainly determined by $T$, that is, the number of users that the decoder tries to recover in each slot. Note that TIN-SIC is more preferable than joint decoding since it utilizes single-user decoding, which makes it generally less complex and more parallelizable. Spreading-based and ODMA-based schemes, proposed in [94] and [78], respectively, have higher complexities compared to their counterparts in exchange for their superior performance. For instance, the complexity of MMSE filtering for fading coefficient estimation in [78] is cubic in $K_{a}$.
V-C MIMO Fading Channels
We now turn our focus to the URA solutions for MIMO fading channels. Several studies have investigated practical coding schemes for the MIMO URA setup, which can be categorized into five groups: CCS-based, tensor-based, pilot-based, ODMA-based, and Bayesian approaches, with inevitable overlaps between them. The MIMO URA solutions based on these approaches are reviewed in the rest of this subsection.
Figure 14: Encoding and decoding structures of tensor-based schemes for $d=2$ [109].
V-C1 CCS-based schemes
In CCS-based algorithms for MIMO fading URA [97, 98], the bit sequence of each user is divided into $S$ sub-messages, which are then mapped to codewords in pilot codebooks with lower dimensions to obtain $S$ different pilots, transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot, and they are then combined to form the complete length-$B$ bit sequence for each user. To establish connections between these sub-messages, the authors in [97] incorporate parity-check bits into each sub-message, while [98] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the necessity for additional parity bits enhances the performance of [98] in comparison to [97].
Eliminating the parity bits in [97] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [99, 102, 100, 101]. To stitch the decoded sub-messages, the authors of [99] employ a joint Bayesian decoding exploiting the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [100], and the similarity of the estimated channel vectors are used in [101]. The structure of the inner SPARC code allows parallel single-user outer decoding of user messages in [102], which mitigates the need for the parity check bits. An uncoupled structure is employed in [103] to design a CCS-based URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [104], where a long sub-block recovered in the first phase is used for channel estimation, and very short lengths of sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [105] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller one in both the codeword length and the number of codewords. Moreover, an uncoupled CS scheme employing polar coding and power diversity in combination with AMP-based decoding and SIC at the receiver side is proposed in [106].
V-C2 Tensor-based schemes
Another trend in MIMO URA is the tensor-based schemes [107, 108, 109], where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor). For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [107], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit-sequence is detected. By employing OFDM transmission, this scheme is extended to the case with timing offsets in [108]. On the other hand, the authors of [109] propose a block-term decomposition for the decomposition of 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. The overall coding structure of tensor-based methods is depicted in Fig. 14.
Figure 15: Encoding structure of pilot-based schemes.
Figure 16: An illustration of the ODMA idea for the fading scenario with the pilot and data parts. Colored boxes show the utilized time instances in the data part.
V-C3 Pilot-based schemes
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems, where, the idea is to divide user bit sequences into pilot and data parts. In the existing solutions, the transmitted signal of each user also consists of two parts, as depicted in Fig. 15: the pilot part which is obtained by mapping the pilot bits to a codeword in a pilot codebook (a non-orthogonal codebook [111, 12, 113, 112, 14, 110], or a codebook consisting of multiple stages of orthogonal pilots [13]), and the data part which is encoded by a channel code (a polar [13, 111, 12, 14] or an LDPC code [113, 112]). For decoding, as shown in Fig. 17, these schemes iteratively employ the following steps:
  • At the initial stage, the receiver detects the active pilots using a user activity detection technique, such as generalized OMP (gOMP) [14, 110], AMP [112, 114], multiple-measurement vector AMP (MMV-AMP) [111], energy detector [12], NP hypothesis testing [13] or CB-MLD [113].
  • The channel vectors corresponding to each detected pilot are estimated by techniques such as MMSE [111, 12, 14, 114, 110], MRC [13] or AMP [113, 112].
  • Using the estimated channel coefficients, soft estimates of the transmitted codewords are obtained via decoupling techniques, such as MRC [111, 13, 14, 110], MMSE [12] or AMP [113], which are then used for generating log-likelihood ratio (LLR) values to be fed to single-user decoders [13, 111, 12, 113, 14, 114] or a joint factor graph [112].
  • In most pilot-based schemes, following the SIC approach, the contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the next iterations. Note that in [13, 12, 14, 110], the channel vectors are re-estimated using the successfully decoded codewords to have a better interference cancellation at the SIC stage.
Figure 17: Decoding structure of pilot-based schemes.
V-C4 ODMA-based schemes
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is extended to the fading scenario as well for the case of a single-antenna receiver in [78], and a massive MIMO receiver in [110, 115, 116]. The main idea in [110] is to allocate a part of the time frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the time frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 16. At the receiver side, the gOMP algorithm is employed to detect the active pilots, followed by LMMSE channel estimation, MRC symbol estimation, single-user polar decoding, and SIC to recover the data part. Note that if the channel is a GMAC, the entire time frame is used for data transmission. On the other hand, the authors of [115] consider superimposed pilot and data transmission, where the pilot sequence occupies the entire time frame and data signal is transmitted by ODMA. The receiver operation starts with an energy detection-aided non-Bayesian pilot detection, and the same steps as in [110] are utilized for the decoding of the data part. A slotted ODMA-based scheme combining preamble transmission and convolutional LDPC coding is proposed in [117], with AMP decoding for the preamble and turbo probabilistic data association for multiuser detection of the data part at the receiver side. Furthermore, two slotted pilot-free ODMA schemes are proposed in [116] and [118]. While an iterative TIN strategy is employed for joint pattern and data detection in [116], random dictionary learning is studied in [118] in conjunction with clustering-aided activity detection.
V-C5 Bayesian approach
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently. Hence, the error from one block is propagated to the next one. To resolve this issue, there is also a new set of MIMO URA schemes that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal, and their corresponding channel coefficients [120, 119]. Specifically, in [119], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Besides, the authors of [120] propose to encode the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
Figure 18: The decoding steps of the main URA approaches for MIMO fading channels.
Summary and Performance Comparison:
A summary of MIMO URA schemes is provided in Table V, accompanied by the decoding steps of the main approaches summarized in Fig. 18. Moreover, in Fig. 19, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients. The results in Fig. 19 show that due to adopting strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes in [111, 12, 13, 14, 110] usually outperform the tensor-based scheme in [107] for which a sub-optimal coding scheme is utilized, while the SKP coding-based scheme in [119], following the Bayesian approach, offers similar performance to the best performing pilot-based schemes. Finally, note that, in order to benefit from increased spatial diversity, a receiver with a massive number of antennas (i.e., a massive MIMO setup) is considered in the provided numerical examples in most MIMO URA papers. However, in order for the suitable URA candidate for next-generation applications to cover a wider range of applications, a comprehensive comparison of the state-of-the-art schemes is required, especially for a smaller number of receiver antennas.
Computational Complexity:
Figure 19: Comparison of URA schemes for MIMO fading channels with a target PUPE of $0.05$, $n=3200$, and $B=100$ (the common set of parameters in the literature).
Table V: A summary of URA solutions for MIMO fading channels.
Paper Approach Description
Gao et al., 2023 [86] - Provides achievability and converse bounds.
Fengler et al., 2021 [97] CCS-based The pilots at each slot are found using a newly proposed coordinate-wise activity detector. Then, the detected pilots of different slots are stitched together using parity bits.
Shyianov et al., 2021 [98] CCS-based Instead of employing parity-check bits (as in [97]), the correlation between different slots is used for stitching pilots, hence, the energy efficiency is improved.
Jiang et al., 2021 [104] CCS-based A two-stage scheme with a long sub-message in the first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage.
Decurninge et al., 2020 [107] tensor-based Users’ signal separation and detection is done by tensor decomposition.
Luan et al., 2022 [109] tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [107].
Fengler et al., 2022 [111] pilot-based A pilot and a polar codeword are transmitted sequentially. For decoding, the active pilots and corresponding channel coefficients are detected using MMV-AMP and MMSE, respectively, and a polar decoder is employed for data detection.
Ahmadi et al., 2024 [13] pilot-based Each user transmits its multiple stages of orthogonal pilots along with a polar codeword through a randomly selected slot. An energy detector, an MRC, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively, followed by an LS-based SIC.
Gkagkos et al., 2023 [12] pilot-based A pilot and a random spread polar codeword are transmitted sequentially. An energy detector, an MMSE, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, a SIC removes the contribution of successfully decoded signals.
Su et al., 2023 [112] pilot-based A pilot is transmitted along with an LDPC coded and modulated codeword. At the receiver side, AMP is used for pilot detection and channel estimation, and an LDPC decoder is used for data detection. The selected pilot specifies the indexes over which every symbol of the LDPC codeword is transmitted.
Su et al., 2023 [113] pilot-based A pilot and an LDPC codeword are spread and transmitted consecutively. At the receiver’s end, a maximum likelihood detection (MLD) algorithm detects active pilots, an AMP technique estimates the corresponding channels, and a SIC block is applied following the LDPC decoder.
Ozates et al., 2024 [14] pilot-based The transmission time frame is slotted, and each user randomly selects a slot to transmit a pilot along with a polar codeword. For decoding, a generalized OMP and an MMSE estimator are used for pilot detection and channel estimation, the polar decoder is used for data detection, followed by SIC.
Ozates et al., 2024 [110] pilot-based The transmission time frame is divided into pilot and data parts, and each user transmits a non-orthogonal pilot followed by its polar codeword based on ODMA. At the receiver, an iterative decoder employing gOMP, LMMSE, MRC, single-user decoding, and SIC is utilized.
Han et al., 2023 [119] Bayesian Employs FEC code, and sparse spreading.
Jiang et al., 2023 [120] Bayesian Combines sparse regression codes for encoding and a message-passing algorithm for decoding.
When comparing the computational complexity of the coding strategies for MIMO URA, CCS exhibits the lowest complexity among them, as its complexity is linear in $M_{r}$ and the number of common codebook elements [97]. Tensor-based [107] and non-slotted pilot-based [12] schemes have a similar complexity that is cubic in $K_{a}$. This complexity results from the canonical polyadic decomposition in the former and MMSE filtering for channel estimation in the latter. On the other hand, the complexity of slotted pilot-based schemes [13], [14] is cubic in the number of active users in the slot, which demonstrates the complexity advantage of slotting. Furthermore, the complexity of the Bayesian approach in [119] is linear in $K_{a}$, $M_{r}$, and $n$. To summarize, for practical system parameters, CCS has the lowest complexity, while the complexities of slotted pilot-based schemes and the Bayesian approach are similar and smaller than those of tensor-based and non-slotted pilot-based schemes.
V-D Existing URA Solutions on Alternative Setups
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution can be employing RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, it is shown that a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of a URA system by enabling line-of-sight paths between the transmitters and the receivers in problematic environments with many blocking obstacles [121, 122]. More specifically, the RIS-based schemes in [123, 124] operate in two phases. In the first phase, called the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, called the data phase, transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [123] outperforms the one in [124], with a power saving of up to 10 dBm. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, as well as a more suitable metric in the RIS design in [123], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [124]. The authors improve the RIS design metrics of [123] in [125] and obtain some bounds on its performance.
Conventionally, in MIMO URA systems, the ideal hardware units are assumed at both the BS and the UE side. However, due to the massive number of users and the antenna elements, the hardware needs to be inexpensive, and these elements are prone to hardware impairments (HWI) such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWI is considered in [126], where the solution developed in [14] is adapted to the presence of HWI.
While most URA works consider the transmission over i.i.d. Rayleigh block-fading channels, some investigate other channel models. A channel model with one single dominant path between a user and the BS is considered in [127], and an approach similar to the one in [114] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [128], namely, the users can start or end their transmission in any slot and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing. Type-based MA, where each device observes the state of a process and transmits it to a common receiver is investigated in an unsourced setup in [129], where the receiver’s goal is to estimate the states and their multiplicity. The problem of user identification and authentication is applied to URA in [130] with the aid of a message authentication code. Moreover, the authors of [131] propose a URA scheme combining physical and medium access layer approaches where medium access layer slotting carries information. They show that carrying information by the medium access control layer improves the system performance. The Rician fading channel is considered in [132] through an application of coded compressed sensing. A correlated fading scenario is tackled in [133]. Frequency-selective fading is examined in the context of MIMO URA in [134] and [135]. The authors of [134] propose a tensor-based modulation scheme, while a slotted scheme employing OFDM is studied in [135].
The asynchronous setup receives increasing attention due to its easier implementation in hardware, and it is studied in MIMO URA in [136, 138, 137]. A slotted concatenated coding scheme is proposed in [136] based on the SPARCs and tree codes in an OFDM system. The authors of [137] tackle both the timing and frequency offsets by proposing an asynchronous URA scheme by employing SPARC, LDPC codes, and multistage orthogonal pilots. Moreover, symbol asynchronicity in URA is investigated in [138].
Furthermore, some works have investigated URA in other communication setups such as cell-free MIMO [139, 140, 141] and erasure [142, 143] channels. For instance, a cell-free setup is considered in [139] where the coverage area is divided into multiple regions each served by a receiver unit. The users in each region employ a common codebook different from other regions to transmit their messages, and the received signals at the receiver units are jointly processed in a central unit. Joint processing is also considered in [140] and [141], where a combination of channel coding and fixed sparse coding is employed in the former, while an uncoupled CCS scheme utilizing coarse location information for trellis stitching is proposed in the latter.
The common assumption in the URA literature is that all the users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, which is called multi-class URA is addressed in [144, 145] in the coded compressed sensing framework. More specifically, a scenario is considered where devices are clustered into two classes with different SNR levels and payload requirements. In the cluster with higher power, devices transmit using a two-layer superposition modulation, while those in the other cluster transmit with the same base constellation as in the first cluster. In order to recover packets at the receiver, a CCS-based approach is employed, where an outer code is used to stitch message fragments together across times and layers.
Now that we have reviewed the existing URA solutions, we provide a summary of different URA setups and the corresponding solution approaches in Fig. 20.
Figure 20: A summary of different URA setups in the literature and their corresponding common approaches.
VI Remarks and Future Research Directions
In this survey, we had a comprehensive look at the URA literature with an emphasis on the state-of-the-art. We grouped the different URA solutions and explained the representative work in each in more detail while clarifying its connection to related works. Moreover, we compared the works in the literature both setup-wise and performance-wise in order to have a clearer view of the existing URA solution.
As evident from this survey, there have been numerous works on URA in recent years. However, there are still some issues, some critical, that need to be investigated further. We will next outline some of these issues and provide several future research directions. Moreover, we highlight potential methods of addressing them.
VI-A Asynchronous URA
In URA, the common assumption is that the users transmit simultaneously; or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of the users and devices in envisioned URA applications such as massive sensor networks, such transmission coordination is very challenging. This makes asynchronous URA an important research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. The reason is that due to the massive number of potential devices, and hence diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous. It is shown that if the asynchronicity is dealt with properly, it can lead to performance improvement even compared to the synchronous case [146].
There have been some attempts to address the asynchronous setup [91, 90, 95, 136, 137, 51], however, with a bounded time horizon, which enable them to convert the problem to one similar to the synchronous case, e.g., by employing a limited decoding time frame. For instance, in [91], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users.
The only work that focuses on a fully asynchronous (i.e., with an unbounded time horizon) URA setup is [147], where the authors consider a window-based decoding for GMAC URA. In this scheme, each user transmits a preamble (from a pool of available ones) for timing acquisition concatenated with a channel-coded payload, and the receiver uses a SIC-based sliding window for recovering the transmitted packets.
URA-specific challenges: The main URA-specific challenge of a fully asynchronous setup for fading URA is the computational complexity, which is much higher than the conventional (non-URA) setup due to the higher number of required preambles. More specifically, in an asynchronous GMAC setup, the preambles only serve to identify the start of the packets at the receiver, so a preamble pool of even a very small size would suffice; while in an asynchronous fading setup, they are also needed to estimate the channel coefficients, which are then used for user separation. In a conventional fading setup, the receiver only searches over the assigned preambles, whose number is equal to the number of active users. However, in a URA setup, due to the lack of coordination between the users and the receiver, the number of available preambles has to be much higher than in the conventional setup, hence much higher computational complexity for user activity detection and channel estimation. This computational complexity significantly increases in the asynchronous setup with a sliding window receiver, as the receiver shifts and correlates all possible preambles through the window to find the start of the packets as well as the user channel estimates, repeating this process with each shift of the sliding window. This highly increases the overall computational complexity, which can even make the solution practically infeasible. One way to reduce the complexity is to increase the moving steps of the sliding window, at the cost of lower overall performance. However, smarter schemes are needed to alleviate this issue.
VI-B Robust URA
Like any other communication system, the performance of a URA system is affected by the presence of imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments where the employed hardware is not as ideal as assumed, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, which is very probable in many URA applications where the devices are actually cheap and low-precision sensors. Only one work in the URA literature has addressed this issue [126], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensation are statistically modeled as additive and multiplicative noises. Therefore, the analysis of different hardware impairments on the performance of existing URA solutions and providing robust solutions for them is an interesting research direction.
URA-specific challenges: Looking at the URA literature, we can observe that the URA schemes are more sensitive than the conventional ones to the choice of design parameters. That is, in order for the URA schemes to have acceptable performance, their design parameters need to be optimized every time a system hyper-parameter, such as active user load or the target PUPE, changes. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyper-parameters. One way is to employ stochastic optimization techniques, e.g., optimizing the expected value of the cost metric (e.g., PUPE) over the statistics of the intended system hyper-parameters in order to obtain the set of design parameters that work well in a wide range of system hyper-parameters.
VI-C Machine Learning-Aided URA
Current solutions look at the URA problem in a single shot. However, the transmission is continuous and spans several frames in many practical applications. In these cases, machine learning (ML) techniques such as reinforcement learning can be used to adapt to the environment and learn suitable transmission strategies. ML techniques can also be used to simplify some blocks in the decoding process, such as user activity detection (similar to the deep learning based solutions in [148]) and channel estimation. For instance, in [149], the authors show that employing a dictionary learning based solution obviates the need for pilots for channel estimation in the fading URA scenarios while maintaining an acceptable performance. URA-specific challenges: Apart from general challenges in implementing ML solutions in communication systems, such as energy and computational constraints and compatibility with legacy systems, there are some that are specific or more pronounced in a URA setup, which is mostly due to the massive number of potential users (up to the order of millions) and their sporadic activity. Training a neural network with an input of very large size (due to the massive number of potential users) with non-stationary characteristics (due to the sporadic activity) is a very challenging task.
It can be vice versa as well: URA techniques can be employed to better realize some ML-based solutions. For instance, in federated learning (FL), the receiver is only interested in the sum of the transmitted signals (a function of neural network weights or their updates), not them separately. The users transmit simultaneously in over-the-air FL [150], which exploits the additive nature of the wireless channel. This, alongside the huge number of potential users in envisioned FL applications, makes URA a suitable framework for the physical layer [151, 152]. Challenges of URA in FL: Note that in distributed function computation applications, such as FL, where the receiver is not interested in the transmitted messages themselves but only in a function of them (e.g., their summation in FL), PUPE is not a suitable metric for the URA implementation, as it is defined on the individual messages. So, new metrics (even higher-level ML-based ones such as accuracy) need to be employed to better capture the system performance.
VI-D Integrated Sensing and Communication in URA
Existing research on URA has focused primarily on the transmission of information from communication users. However, with the evolution of 6G and beyond, new functionalities such as sensing, computing, and security are becoming critical alongside communication. As the hardware architectures and signal processing methods for these tasks are mostly similar, jointly adopting them significantly reduces hardware costs compared to separate systems while maintaining comparable performance [153]. A promising example of such joint designs is integrated sensing and communication (ISAC) systems.
In [154], the problem of integrating sensing and communication is considered in a URA setup, where a massive number of unsourced and sporadically active users are connected to the system. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievable bound for the UNISAC to show its efficacy in sensing the sensing users as well as in communicating with the communication ones in a URA setup.
There are still many aspects of an ISAC system in a URA setup that need further investigation. For instance, the problem becomes more challenging when sensing users are passive, as opposed to the active assumption in [154]. In this case, only communication users transmit signals to the BS, with the same signal opportunistically used to sense passive targets.
URA-specific challenges: In a URA system, a massive number of users transmit sporadically without being registered or authenticated by the network. Therefore, there is no control over transmitter characteristics such as location, signal structure, and beam direction, which are critical for reliable passive target sensing. Additionally, the random access nature of user transmissions causes heavy interference, a phenomenon uncommon in traditional sensing techniques, resulting in a lack of solutions to mitigate it. Therefore, advanced algorithms are needed to address these challenges and enhance the system’s ability to sense passive targets effectively.
VI-E URA with Feedback
In communication systems, feedback has been shown to greatly simplify transmission and reception mechanisms and, in some cases, even improve system performance [155]. The main challenge in employing feedback-based techniques in URA is the lack of coordination between the users and the receiver due to the massive number of potential users. However, even in URA, there can be some level of coordination as long as the signal sent from the receiver is not user-specific, as there is no user identity in URA. For instance, the receiver can broadcast a common pilot/beacon to the users so that they check if they are in the set of successfully decoded users in the previous round, and estimate their channel coefficients or at least their large-scale components depending on their dedicated time slots. Just note that the time dedicated to this operation needs to be considered as extra overhead when comparing with the solutions without feedback. This issue is partly tackled in [156, 157, 158]. In [156], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques; while in [157], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a pre-defined threshold. One application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [158] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
Recall that the main novelty of URA is providing a unified framework to decouple user identification and data transmission problems. This makes it a fit solution in applications where user identification is not feasible at all due to the massive number of potential users. However, URA is still applicable and a good candidate in non-massive (but with still a considerable number of potential users) scenarios where user identification is still feasible. In these scenarios, conventional authentication and acknowledgment schemes impose a huge overhead on the system. This issue is tackled in [159].
VI-F RIS-Aided URA
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play an important role in improving the overall system performance by providing a substitute communication link with potentially many degrees of freedom. However, the huge channel estimation overhead is the main challenge in the use of RIS in URA systems. The channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the high number of RIS elements and the potentially massive number of antennas at the receiver (as the benefits of a RIS system are amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even users, which renders most existing channel estimation schemes in RIS infeasible. There are a few works that have considered the RIS-aided URA scenario [123, 124, 125]. Although these works make the channel estimation feasible in a RIS-aided URA setup by employing an SIC-based iterative approach and time frame slotting strategies, the high computational complexity still remains a challenge.
Moreover, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [160] and metasurface-based RIS [161], in order to find suitable candidates. Moreover, the employment of massive arrays such as extra large massive MIMO and RIS leads to some active users being in the near field of these arrays, which makes hybrid near field-far field analysis a necessity. This is in contrast with the existing URA solutions, which are all in the far-field regime.
VI-G Standardization
There has not been any effort in the standardization of URA, neither for incorporating it in the next updates of the current protocol nor for future generations such as 6G. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that preferably can support both the current protocols, perhaps with small modifications, and the high demands of the future ones. For instance, in [162], with a small modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with a small number of active users (10 to 50 users), employing an IDMA-based URA solution. However, this solution is not scalable to higher active user loads.
URA-specific challenges: There are two main challenges in incorporating URA in current protocols, such as 5G-NR and beyond, for reasonable active user loads (e.g., in the order of hundreds). The first is the small number of preamble sequences available in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second is the limited number of access patterns that can be realized, which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This shows that an important step in the realization of URA in future protocols is to considerably increase the available pilot or transmission patterns.